### v3

# Asymptotically Euclidean Solutions of the Constraint Equations with Prescribed Asymptotics

Abstract
We demonstrate that in constructing asymptotically flat vacuum initial data sets in General Relativity via the conformal method, certain asymptotic structures may be prescribed a priori through the specified seed data, including the ADM momentum components, the leading- and next-to-leading-order decay rates, and the anisotropy in the metric’s mass term, yielding a recipe to construct initial data sets with desired asymptotics. We numerically construct a simple explicit example of an initial data set, with stronger asymptotics than have been obtained in previous work, such that the evolution of this initial data set does not exhibit the conjectured antipodal symmetry between future and past null infinity.
1 Introduction
[…]
The first work which established the broad existence of solutions to the initial value problem for Einstein’s equation admitting a complete notion of infinity, outside of the few known explicit spacetimes, was Christodoulou and Klainerman’s proof of the nonlinear stability of Minkowski space sta . They proved the stability of key geometric features of the maximal vacuum Cauchy development of sufficiently “small” initial data sets that are asymptotically flat in the sense that there exists a coordinate system $(x_{1},x_{2},x_{3})$ on the complement of a compact set in $\Sigma$ in which [^1] For $f\in C^{m}(\mathbb{R}^{n})$, we define $f=o_{m}(r^{\delta})$ provided that $D^{\alpha}f=o(r^{\delta-|\alpha|})$ as $r\to\infty$ for any multi-index $\alpha$ of order $|\alpha|\leq m$.
$\tilde{g}_{ij}$ $=\left(1+\frac{2M}{r}\right)\delta_{ij}+o_{4}(r^{-3/2}),$ (1.1)
$\widetilde{K}_{ij}$ $=o_{3}(r^{-5/2}).$ (1.2)
We define these falloff conditions as type (CK). This conclusion was later generalized by Bieri lydia1 ; lydia2 to initial data sets with the relaxed falloff conditions (type (B))
$\tilde{g}_{ij}$ $=\delta_{ij}+o_{3}(r^{-1/2}),$ (1.3)
$\widetilde{K}_{ij}$ $=o_{2}(r^{-3/2}).$ (1.4)
In each case, one has different degrees of control on the induced falloff rates of various curvature components as one approaches $\mathscr{I}^{\pm}$. As also investigated by Bieri lydia4 ; lydia14 ; lydia5 , different control still is afforded by the intermediate falloff conditions (type (A))
$\tilde{g}_{ij}$ $=\delta_{ij}+h_{ij}+o_{3}(r^{-3/2}),$ (1.5)
$\widetilde{K}_{ij}$ $=o_{2}(r^{-5/2}),$ (1.6)
where each component of $h_{ij}$ is homogeneous of degree 1-1, representing an anisotropic mass.
To better understand the asymptotic behavior of the geometry of spacetimes compatible with these three falloff conditions, and to assess the status of physically-motivated conjectures associated to that geometry, it is interesting and necessary to construct examples of constraint-satisfying initial data sets in each of the classes (CK), (B), and (A), requiring the ability to prescribe the data’s falloff rate. While the ultimate act of construction is a numerical task, it is vital to first develop a mathematical blueprint of the procedure underpinned by theorems guaranteeing its efficacy. A subset of the present authors bieri_brill_2025 recently provided such a blueprint in the rather restrictive setting of Brill wave initial data sets, requiring both azimuthal and time symmetry. These results, however, were insufficient to guarantee the construction of the more delicate type (A) data. That work numerically constructed an initial data set of type (B) whose evolution cannot satisfy Strominger’s antipodal conjecture in strominger , which is a topic of interest in the literature symmprabhuetal ; symmmagdyetal ; lz1 . That conjecture posits a symmetry of a particular component of the electric part of the Weyl curvature ($\Psi_{2}$ in the Newman-Penrose formalism, or $\rho$ in Christodoulou and Klainerman’s notation) evaluated asymptotically along $\mathscr{I}^{+}$ as compared to $\mathscr{I}^{-}$: The conjecture says that if one evaluates this component’s limit on each of $\mathscr{I}^{\pm}$ and proceeds along each to spatial infinity, the two resulting functions on the sphere should agree up to composition with the antipodal map on the sphere, $p\mapsto-p$.
In this work, we significantly generalize the results of bieri_brill_2025 both to do away with any symmetry assumptions and to handle type (A) data, establishing a sequence of results guaranteeing that one can control the falloff rates and the ADM momenta of initial data sets generated via the widely utilized conformal method. We use these results to construct an initial data set of type (A) whose evolution cannot satisfy the antipodal conjecture of strominger , restricting the class of spacetimes to which this conjecture could apply. Since the conjecture holds trivially for spacetimes evolved from type (CK) data, our results indicate that the domain on which the conjecture might hold nontrivial content must be somewhat small. For simplicity of presentation, we work with $\Sigma=\mathbb{R}^{n}$, but the analytical results generalize straightforwardly to all asymptotically Euclidean manifolds, even with multiple ends. Of course, one takes $n=3$ in standard general relativity (as above), but we have found it interesting and instructive to generalize the mathematical results presented here to arbitrary dimension $n\geq 3$.
We work with finite regularity measured in Sobolev scales, and in particular with weighted Sobolev spaces. This common setting for building asymptotically Euclidean initial data usually focuses on decay rates within the isomorphism range of the applicable elliptic operators (e.g. bruhat-isenberg , maxwell_rough_2006 ). While this approach suffices for establishing existence and uniqueness, the slow decay rates lose asymptotic information. We work instead with faster decay rates that expose asymptotic structure, and with particular care at the critical transition rates where the Fredholm index of the operators changes (see Proposition 3.6 and Lemma 4.6). Although we work exclusively in the context of constant mean curvature (CMC) solutions of the constraints, the tools developed here are equally applicable to perturbative constructions of non-CMC initial data.
We remark that Dain and Friedrich published a work in 2001 dain2001asymptotically having a fair bit of conceptual overlap with (and a remarkably similar title to) the present work. They also proved a number of interesting results allowing control of asymptotic features of initial data sets constructed via the conformal method, including the prescription of a leading-order Euclidean transverse traceless tensor in $\widetilde{K}_{ij}$ allowing control of the ADM momentum (cf. our Theorem 5.3). However, there is little technical overlap with the present work, either in precisely what is proved or in the analytical techniques employed. Namely, Dain and Friedrich worked (appropriately to their objectives) in the restricted setting of initial data sets admitting a conformal compactification at spatial infinity, so they were interested in obtaining asymptotics which amount to analyticity at spatial infinity, asking that $\tilde{g}_{ij}$ and $\widetilde{K}_{ij}$ admit a full power series expansion in integer powers of $1/r$. The present work seeks to provide a simple procedure for constructing initial data sets in a much broader category, with emphasis on controlling the structure of the leading-order terms (especially in $\tilde{g}_{ij}$) and the precise (non-integral) decay rate of subleading terms, and making no demand of a full power series expansion. Broadly, we are interested in the construction of initial data sets conforming to varying notions of asymptotic flatness for the ultimate purpose of probing the extent to which physically interesting features of spacetime depend on this notion, but the results in dain2001asymptotically are not amenable to handling a variety of falloff behaviors. In particular, results in dain2001asymptotically could be described as allowing the construction of highly specialized examples of type (CK) data, but they are not relevant to the construction of type (A) data which is not (CK), which is among our chief concerns.
In a similar vein, in 1996 Beig and O’Murchadha used spherical harmonic expansions to investigate the asymptotics of solutions to the CMC vacuum momentum constraint equation (2.4) beig_momentum_1996 . Working in the smooth setting in dimension $n=3$, they fully characterized the asymptotic forms and decay rates of solutions to the momentum constraint with seed data having given arbitrarily fast, non-exceptional falloff compatible with type (CK) data. This characterization was made up to a collection of constant “multipole moments”, including the ADM momentum components, which can be prescribed via an appropriate choice of seed data (but not anticipated for an arbitrary choice of seed data). We have built upon Beig and O’Murchadha’s work by doing the following: extending some of their conclusions to lower regularity and arbitrary dimension; establishing that similar conclusions hold for solutions to the Hamiltonian constraint equation (2.3) (and hence one’s entire initial data set); showing that the ADM momentum components can be computed directly from the seed data; and investigating the asymptotics of solutions if the seed data has the exceptional “threshold” falloff $O(r^{2-n})$, allowing the construction of type (A) data which is not (CK).
[…]
2 Preliminaries and The Conformal Method
[…]
$R(\tilde{g})-|\widetilde{K}|_{\tilde{g}}^{2}+\text{tr}_{\tilde{g}}(\widetilde{K})^{2}$ $=0,$ (2.1)
$)\displaystyle\text{div}_{\tilde{g}}(\widetilde{K})-d(\text{tr}_{\tilde{g}}(\widetilde{K}$)) $=0.$ (2.2)
The first of these is the Hamiltonian constraint, and the second is the momentum constraint. There is a broad literature on the study of the constraint equations, spanning from methods of constructing solutions, either from scratch or from existing solutions, to properties that solutions enjoy, such as the positive mass theorem and the Riemannian Penrose inequality. We are interested in the former, and we seek to construct initial data sets corresponding to a maximal time slice in $\widehat{M}$, wherein $\text{tr}\,_{\tilde{g}}(\widetilde{K})=0$. The vacuum constraint equations then take the following form:
[…]
We pose these equations on $\mathbb{R}^{n}$, to be solved for $\tilde{g}$ and $\widetilde{K}$. These are underdetermined, as they impose only $n+2$ conditions (including $\text{tr}\,_{\tilde{g}}(\widetilde{K})=0$) on the two symmetric (0,2)-tensors $\tilde{g}$ and $\widetilde{K}$. To determine a parameterized set of solutions, one must supplement these conditions, effectively specifying part of $\tilde{g}$ and $\widetilde{K}$ and leaving the rest of their content to be determined by the constraints.
A historically powerful approach in the analysis of these equations, known as the conformal method bruhat-york ; lichnerowicz1944 , is to specify the conformal class of $\tilde{g}$, amounting to stipulating an arbitrary “unphysical” metric $g$, as well as a traceless and symmetric $(0,2)$-tensor $A_{ab}$ (the conformal velocity) and a scalar function $N=1+\delta N:\mathbb{R}^{n}\to\mathbb{R}_{+}$ (the lapse function). Once this seed data comprised of $g$, $A_{ab}$, and $\delta N$ has been specified on $\mathbb{R}^{n}$, one sets
$K_{ab}:=\frac{1}{2N}\left(A_{ab}-(\mathbb{L}_{g}W)_{ab}\right)$
and $q_{n}:=\frac{2n}{n-2}$ and proceeds to solve the PDE system
$-(q_{n}+2)\Delta_{g}\varphi+R(g)\varphi$ $=\varphi^{-q_{n}-1}|K|_{g}^{2},$ (2.5)
$\text{div}_{g}\left(\frac{1}{2N}(\mathbb{L}_{g}W)_{ab}\right)$ $=\text{div}_{g}\left(\frac{1}{2N}A_{ab}\right)$ (2.6)
for the vector field $W^{a}$ and the positive scalar field $\varphi$. Here, $\mathbb{L}_{g}$ is the conformal Killing operator, which is defined to be the following symmetrized and traceless covariant derivative:
$\mathbb{(}\mathbb{L}_{g}W)_{ab}:=\nabla_{a}W_{b}+\nabla_{b}W_{a}-\frac{2}{n}g_{ab}\nabla_{c}W^{c}.$
Once $W^{a}$ and $\varphi$ have been obtained by solving equations (2.5) and (2.6), one sets
$\tilde{g}$ $:=\varphi^{q_{n}-2}g,$
$\widetilde{K}_{ab}$ $:=\varphi^{-2}K_{ab}.$
Using the conformal divergence identity $\text{div}_{\tilde{g}}(\varphi^{-2}K)=\varphi^{-q_{n}}\text{div}_{g}(K)$ for a traceless and symmetric $(0,2)$-tensor $K_{ab}$ as well as the well-known conformal transformation of scalar curvature, one readily confirms that this choice of the physical metric $\tilde{g}$ and the second fundamental form $\widetilde{K}_{ab}$ solves the constraint equations (2.3)-(2.4).
A key feature of the conformal method for maximal data is that the conformal constraint equations (2.5)-(2.6) are semi-decoupled in the sense that (2.6) can be solved for $W^{a}$ independently of $\varphi$, and then equation (2.5)– the Lichnerowicz equation– can be solved for $\varphi$. We note that in the Einstein vacuum case discussed here, these equations admit unique solutions for each choice of the seed data in appropriate function spaces (discussed below) maxwell_solutions_2005 .
We are interested in prescribing the falloff rates of $\tilde{g}$ and $\widetilde{K}$ by means of building them into the seed data $g$, $A_{ab}$, and $\delta N$. To leverage analytical control over the decay rates of these quantities, we operate in weighted Sobolev spaces, which we now define:
Definition 2.1.
Let $k\in\mathbb{Z}_{\geq 0}$, $1<p<\infty$ and $\delta\in\mathbb{R}$. The weighted Sobolev space $W^{k,p}_{\delta}$ consists of the functions $u:\mathbb{R}^{n}\to\mathbb{R}$ such that
$u||_{W^{k,p}_{\delta}}^{p}:=\sum_{|\alpha|\leq k}\int_{\mathbb{R}^{n}}\left<x\right>^{-n-p\delta}|\partial_{\alpha}u|^{p}\;dx<\infty,$
where $\left<x\right>=\sqrt{1+|x|^{2}}$. The analogous norm for tensors on $\mathbb{R}^{n}$ is defined component-wise, and we use the same notation $W^{k,p}_{\delta}$ for spaces of tensor-valued functions, leaving the tensor bundle implicit. We also introduce the following notation:
$W^{k,p}_{\delta^{+}}:=\bigcap_{\epsilon>0}W^{k,p}_{\delta+\epsilon},\qquad\qquad W^{k,p}_{\delta^{-}}:=\bigcup_{\epsilon>0}W^{k,p}_{\delta-\epsilon}$
[…]
Definition 2.2.
A metric $g_{ab}\in W^{k,p}_{\mathrm{loc}}$ with $1<p<\infty$ and $k\in\mathbb{N}$ satisfying $k>n/p$ is asymptotically Euclidean of class $W^{k,p}_{\tau}$ for some $\tau<0$ if
$g_{ab}-\delta_{ab}\in W^{k,p}_{\tau}.$
We note that the condition $k>n/p$ ensures that $g_{ab}$ is Hölder continuous and, since $\tau<0$, that it converges uniformly to $\delta_{ab}$ as $x\to\infty$. In keeping with this definition, we work with metrics and lapses in the following category:
Assumption 2.3.
The metric $g_{ab}$ and the lapse $N$ satisfy the following two conditions for some $\tau<0$, $k\in\mathbb{N}$, and $1<p<\infty$ with $k>1+n/p$:
  • $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{\tau}$.
  • NN is a positive function with $\delta N=N-1\in W^{k,p}_{\tau}$
We remark that the core results of this work are true under the relaxed hypothesis $k>n/p$, provided that one additionally assumes that $g$ admits no nontrivial conformal Killing fields (vector fields in the kernel of $\operatorname{{\mathbb{L}}}$) of class $W^{k,p}_{\delta}$ for any $\delta<0$, i.e. which vanish at infinity [^2] It is a reasonable conjecture that this additional assumption is implied by $k>n/p$, but a proof remains elusive.. The following essential multiplication lemma is an immediate consequence of the more general Lemma A.3 discussed in Appendix A, and we note that it suffices for all of our work here.
Lemma 2.4.
Suppose $1<p<\infty$, $k_{1},k_{2},j\in\mathbb{Z}_{\geq 0}$ and $\delta_{1},\delta_{2}\in\mathbb{R}$. Pointwise multiplication of functions in $C^{\infty}_{c}(\mathbb{R}^{n})$ extends to a continuous bilinear map $W^{k_{1},p}_{\delta_{1}}\times W^{k_{2},p}_{\delta_{2}}\rightarrow W^{j,p}_{\delta_{1}+\delta_{2}}$ if
$j$ $\leq\min(k_{1},k_{2}),$
$j$ $<k_{1}+k_{2}-\frac{n}{p}.$
In particular, we may choose the optimal value $j=\min(k_{1},k_{2})$ so long as $\max(k_{1},k_{2})>n/p$.
3 The Momentum Constraint
The asymptotics of a second fundamental form constructed using the conformal method follow from the mapping properties of the lapse-weighted vector Laplacian $P_{g,N}:=\operatorname{\mathrm{div}}_{g}(\frac{1}{2N}\operatorname{{\mathbb{L}}})$ appearing in equation (2.6) along with the asymptotic structure of its source term, the conformal velocity $A_{ab}$. Indeed, setting $Z_{a}:=\nabla^{b}\left(\frac{1}{2N}A_{ab}\right)$, equation (2.6) becomes the inhomogenenous linear equation
$(P_{g,N}W)_{a}=Z_{a}$ (3.1)
which is the main focus of this section.
A routine computation using Lemma 2.4, under Assumption 2.3, yields that for a smooth vector field $X^{a}$ one has
$(P_{g,N}X)_{a}=(\overline{P}X)_{a}+\sum_{|\alpha|\leq 2}B_{ab}^{\alpha}\partial_{\alpha}X^{b},$ (3.2)
where $\overline{P}:=P_{\overline{g},1}=\frac{1}{2}\overline{\operatorname{\mathrm{div}}}(\overline{\operatorname{{\mathbb{L}}}})$ is the Euclidean vector Laplacian (note that we have an extra factor of $1/2$ relative to some other works), associated with the Euclidean metric $\overline{g}_{ab}=\delta_{ab}$, and where each $B_{ab}^{\alpha}\in W^{k-2+|\alpha|,p}_{\tau-2+|\alpha|}$ depends on $g$ and $N$. Since $\tau-2+|\alpha|<0$ for each of these coefficents, the mapping properties of $P_{g,N}$ can be deduced from those of $\overline{P}$. Appendix A contains these details, which imply the following:
Proposition 3.1.
Suppose that $g_{ab}$ and $N$ satisfy Assumption 2.3 with parameters $k$, $p$ and $\tau$. Then for any $\delta\in\mathbb{R}$, $P_{g,N}:W^{k,p}_{\delta}\to W^{k-2,p}_{\delta-2}$ is continuous. If $\delta$ is non-exceptional, then
  1. (a)
    $P_{g,N}$ acting between these spaces is Fredholm,
  2. (b)
    its Fredholm index equals that of the Euclidean vector Laplacian $\overline{P}$ acting between the same spaces,
  3. (c)
    if $\delta<0$ then its kernel is trivial,
  4. (d)
    if $\delta>2-n$ then it is surjective,
  5. (e)
    if $2-n<\delta<0$ then it is an isomorphism,
  6. (f)
    given $V_{a}\in W^{k-2,p}_{\delta-2}$, the equation
    $(P_{g,N}X)_{a}=V_{a}$
    is solvable for $X^{a}\in W^{k,p}_{\delta}$ if and only if $\int_{\mathbb{R}^{n}}V^{a}k_{a}\,dV_{g}=0$ for all vector fields $k^{a}$ in the kernel of $P_{g,N}$ acting on $W^{k,p}_{2-n-\delta}$.
Proof.
The continuity of $P_{g,N}$ follows from Lemma 2.4, and parts (a), (b), and (f) follow from Proposition A.15. Since $k>n/p+1$, Theorem 6.4 of maxwell_solutions_2005 implies that $g_{ab}$ admits no nontrivial conformal Killing fields in $W^{k,p}_{\delta}$ for any $\delta<0$, which is equivalent to part (c). Part (d) now follows from parts (c) and (f), and part (e) is immediate from parts (c) and (d).
The isomorphism range $2-n<\delta<0$ appearing Proposition 3.1 is well studied in the literature and is a routine tool used to find solutions of the momentum constraint (e.g. bruhat-isenberg , maxwell_rough_2006 ). Understanding the asymptotics of these solutions, however, requires a finer analysis of the next decay range $1-n<\delta<2-n$. Part (f) relates solvability in this range to elements of the kernel of $P_{g,N}$ that are bounded at infinity, and we have the following characterization of this kernel:
Proposition 3.2.
Suppose that $g_{ab}$ and $N$ satisfy Assumption 2.3 with parameters $k$, $p$ and $\tau$. If $0<\delta<1$, then the kernel of $P_{g,N}$ acting on $W^{k,p}_{\delta}$ has dimension $n$. Moreover, there is a basis for this kernel consisting of vector fields $k_{(i)}^{a}$, $i=1,\ldots,n$ satisfying [^3] We use parenthetical indices to denote simple labels, distinguished from the usual tensorial indices.
$k_{(i)}^{a}-e_{(i)}^{a}\in W^{k,p}_{\tau}\cup W^{k,p}_{(2-n)^{+}}$
where $e_{(i)}$ denotes the $i^{\text{th}}$ standard coordinate basis vector on $\mathbb{R}^{n}$.
Proof.
Proposition 3.1 implies that the maps $P_{g,N}$ and $\overline{P}$ acting on $W^{k,p}_{\delta}$ are both Fredholm with the same index, and that they are both surjective (since $\delta>2-n$). The dimensions of their kernels therefore agree. Since Proposition A.8 indicates that the kernel of $\overline{P}$ consists only of polynomials, which must be constant since $0<\delta<1$, it follows that $\{e_{(i)}^{a}\}_{i=1}^{n}$ comprises a basis of this kernel, and it has dimension $n$.
Equation (3.2) and Lemma 2.4 yield
$(P_{g,N}e_{(i)})_{a}=((P_{g,N}-\overline{P})e_{(i)})_{a}\in W^{k-2,p}_{\tau-2}.$
We pick $\eta$ satisfying $\eta\geq\tau$ and $2-n<\eta<0$. Since $W^{k-2,p}_{\tau-2}\subset W^{k-2,p}_{\eta-2}$, and since $P_{g,N}$ acting on $W^{k,p}_{\eta}$ is an isomorphism, we can find $U_{(i)}^{a}\in W^{k,p}_{\eta}$ satisfying
$(P_{g,N}U_{(i)})_{a}=-(P_{g,N}e_{(i)})_{a}.$
Uniqueness of the solution implies that the choice of $\eta$ is inessential and $U_{(i)}^{a}\in W^{k,p}_{\tau}\cup W^{k,p}_{(2-n)^{+}}$.
By construction $k_{(i)}^{a}:=e_{(i)}^{a}+U_{(i)}^{a}\in\ker P_{g,N}$. Moreover, since the perturbations $U_{(i)}^{a}$ vanish at infinity, it follows that the $n$ vector fields $k_{(i)}^{a}$ are linearly independent, forming a basis for $\ker P_{g,N}$.
Now suppose $Z^{a}\in W^{k-2,p}_{\delta-2}$ with fast decay $1-n<\delta<2-n$. Although Proposition 3.1 ensures equation (3.1) can be solved for $W^{a}\in W^{k,p}_{(2-n)^{+}}$, the solution does not have faster decay in general. Instead, Proposition 3.1(f) indicates that $W^{a}\in W^{k,p}_{\delta}$ only if each of the integrals
$\mathcal{O}_{i}(Z_{a}):=\int_{\mathbb{R}^{n}}Z_{a}k_{(i)}^{a}dV_{g}$ (3.3)
vanishes. In fact, these obstructions turn out to be closely related to the ADM momentum, which we now examine.
Consider a $C^{1}$ symmetric, traceless tensor $S_{ab}$ with $O(r^{1-n})$ growth. Assuming units with $8\pi G=1$ the ADM momentum of $S_{ab}$ has components
$\mathcal{P}_{i}(S_{ab}):=\lim_{R\to\infty}\int_{\partial B_{R}}S_{ab}n^{a}e_{(i)}^{b}\;d\overline{A},$ (3.4)
with $n^{a}:=x^{a}/r$ the Euclidean radial unit vector, so long as the limit exists. We can express this quantity equally well in terms of the metric $g_{ab}$, its area element $dA$, its normal vector $\nu^{a}$ to the spheres, and the kernel basis $k^{a}_{(i)}$. Because $n^{a}-\nu^{a}$, $dA-d\overline{A}$ and $k^{a}_{(i)}-e^{a}_{(i)}$ each decays at the rate $O(r^{\tau})$, we find
$\mathcal{P}_{i}(S_{ab}):=\lim_{R\to\infty}\int_{\partial B_{R}}S_{ab}\nu^{a}k_{(i)}^{b}\;dA.$ (3.5)
The divergence theorem then yields
$\int_{\partial B_{R}}S_{ab}\nu^{a}k_{(i)}^{b}\;dA$ $=\int_{B_{R}}\nabla^{a}(S_{ab}k_{(i)}^{b})\;dA$
$=\int_{B_{R}}S_{ab}(\nabla^{a}k_{(i)}^{b})\;dA+\int_{B_{R}}(\nabla^{a}S_{ab})k_{(i)}^{b}\;dA.$
If $(\nabla^{a}S_{ab})k_{(i)}^{b}$ and $S_{ab}(\nabla^{a}k_{(i)}^{b})$ are $L^{1}$ functions, we can take a limit in $R$, and we define
$\mathcal{G}_{i}(S_{ab})$ $:=\int_{\mathbb{R}^{n}}S_{ab}\nabla^{a}k_{(i)}^{b}dV_{g},$
$\mathcal{R}_{i}(S_{ab})$ $:=\int_{\mathbb{R}^{n}}(\nabla^{a}S_{ab})k_{(i)}^{b}dV_{g}.$
Note that the quantity $\mathcal{R}_{i}$ defined here for symmetric, trace-free $(0,2)$-tensors is related to the obstruction values $\mathcal{O}_{i}$ for covectors by
$\mathcal{R}_{i}(S_{ab})=\mathcal{O}_{i}(\nabla^{a}S_{ab}).$
The integrability conditions for the integrands defining $\mathcal{G}_{i}(S_{ab})$ and $\mathcal{R}_{i}(S_{ab})$ ensure that the ADM momentum $\mathcal{P}_{i}$ is well defined if both $\mathcal{G}_{i}$ and $\mathcal{R}_{i}$ are, and we have
$\mathcal{P}_{i}(S_{ab})=\mathcal{G}_{i}(S_{ab})+\mathcal{R}_{i}(S_{ab}).$ (3.6)
This decomposition depends on the specific kernel basis $k^{a}_{(i)}$. In particular, since the operator $P_{g,N}$ depends on the lapse, so does the splitting. For reasons we explain below, we call $\mathcal{G}_{i}$ the gravitational momentum, whereas $\mathcal{R}_{i}$ is the residual momentum. For the moment, it suffices to observe that if $S_{ab}$ is trace-free and satisfies the vacuum momentum constraint, then $\mathcal{R}_{i}(S_{ab})=0$ and hence $\mathcal{P}_{i}(S_{ab})=\mathcal{G}_{i}(S_{ab})$.
Now recall the vector field $W^{a}\in W^{k,p}_{(2-n)^{+}}$ introduced above solving $(P_{g,N}W)_{a}=Z_{a}$, where $Z^{a}\in W^{k-2,p}_{\delta-2}$ with $1-n<\delta<2-n$. We claim that the ADM momentum of $1/(2N)(\operatorname{{\mathbb{L}}}W)_{ab}$ is well defined and is determined precisely in terms of the obstruction coefficients of $Z^{a}$:
$\mathcal{P}_{i}\left(\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)_{ab}\right)=\mathcal{O}_{i}(Z_{b}).$
This identity is a consequence of the following lemma, which shows that tensors of the form $1/(2N)(\operatorname{{\mathbb{L}}}W)_{ab}$ cannot carry gravitational momentum in this gauge: that is, $\mathcal{G}_{i}(1/(2N)(\operatorname{{\mathbb{L}}}W)_{ab})=0$. Hence
$\mathcal{P}_{i}\left(\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)_{ab}\right)=\underbrace{\mathcal{G}_{i}\left(\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)_{ab}\right)}_{=0}+\mathcal{R}_{i}\left(\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)_{ab}\right)=\mathcal{O}_{i}(Z_{b}),$ (3.7)
as claimed.
Lemma 3.3.
Suppose that $g_{ab}$ and $N$ satisfy Assumption 2.3 with parameters $k$, $p$ and $\tau$. If $S_{ab}\in W^{k-1,p}_{(1-n)^{+}}$ then $S_{ab}\nabla^{a}k^{b}_{(i)}\in L^{1}$ for each $i$, so that $\mathcal{G}_{i}(S_{ab})$ is well defined. Moreover, if $W^{a}\in W^{k,p}_{(2-n)^{+}}$ then
$\mathcal{G}_{i}\left(\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)_{ab}\right)=0.$ (3.8)
Proof.
We first observe that
$\nabla^{a}k^{b}_{(i)}=\nabla^{a}(k^{b}_{(i)}-e^{b}_{(i)})+\nabla^{a}e^{b}_{(i)}\in W^{k-1,p}_{(-1)^{-}},$ (3.9)
so that Lemma 2.4 implies that $S_{ab}\nabla^{a}k^{b}_{(j)}\in W^{k-1,p}_{(-n)^{-}}\subset L^{1}$ (the condition k1>n/pk-1>n/p ensures that such functions are continuous). Taking $S_{ab}=\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)_{ab}$, we have
$\mathcal{G}_{i}(S_{ab})$ $=\int_{\mathbb{R}^{n}}\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)_{ab}\nabla^{a}k_{(i)}^{b}\,dV_{g}=\int_{\mathbb{R}^{n}}\frac{1}{2N}(\nabla^{a}W^{b})(\operatorname{{\mathbb{L}}}k_{(i)})_{ab}\,dV_{g}$
$=\int_{\mathbb{R}^{n}}\nabla^{a}\left(\frac{1}{2N}W^{b}(\operatorname{{\mathbb{L}}}k_{(i)})_{ab}\right)\,dV_{g}-\int_{\mathbb{R}^{n}}W^{b}(P_{g,N}k_{(i)})_{b}\,dV_{g}$
$=\lim_{R\to\infty}\int_{\partial B_{R}}\frac{1}{2N}(\operatorname{{\mathbb{L}}}k_{(i)})_{ab}W^{b}\nu^{a}\,dA$
since $P_{g,N}k_{(i)}=0$. By equation (3.9) and Lemma 2.4, however, this final integrand is contained in $W^{k-1,p}_{(1-n)^{-}}$, so that the integral vanishes in the limit and $\mathcal{G}_{i}(S_{ab})=0$.
Equation (3.3) suggests a strategy for precisely describing the asymptotics of $W^{a}$ solving equation (3.1) if $Z_{a}$ is rapidly decreasing. We first identify a representative basis $W^{a}_{(i)}$, $i=1,\ldots,n$ of $O(r^{1-n})$ vectors that carry linear momentum: $\mathcal{P}_{i}(W^{a}_{(j)})=\delta_{ij}$. With these in hand, the obstructions $\mathcal{O}_{i}(Z_{a})$ can be removed, and Proposition 3.1(f) then permits finding the zero momentum correction to the solution. The following proposition specifies a construction of such a representative basis, and it yields explicit expressions for the leading order asymptotic expansion of the basis in terms of the Green’s function of the flat vector Laplacian $\overline{P}$,
$G^{ab}=-2C_{n}r^{2-n}(A_{n}\delta^{ab}+B_{n}n^{a}n^{b}),$ (3.10)
with dimensional constants
$A_{n}=3n-2,\qquad B_{n}=(n-2)^{2},\qquad C_{n}=\frac{1}{4(n-1)(n-2)|S^{n-1}|}.$
We remark that we have an explicit factor of $2$ in equation (3.10) due to the factor of $1/2$ in our definition of $\overline{P}$. This Green’s function can often be most effectively leveraged in terms of the $n$ column vector fields $G_{(j)}^{a}$ with components $G_{(j)}^{i}=G^{ij}$. In the statement below, and throughout the remainder of this work, we denote the region exterior to a ball by $E_{R}:=\{x\in\mathbb{R}^{n}\,:\,|x|>R\}$.
Proposition 3.4.
Suppose that $g_{ab}$ and $N$ satisfy Assumption 2.3 with parameters $k$, $p$, and $\tau$. There exist vector fields $W_{(j)}^{a}$, $j=1,\ldots,n$ satisfying the following:
  1. (a)
    $W_{(j)}^{a}\in W^{k,p}_{(2-n)^{+}}$,
  2. (b)
    $(P_{g,N}W_{(j)})_{a}$ is smooth and compactly supported,
  3. (c)
    $\mathcal{P}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W_{(j)}\right)=\mathcal{R}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W_{(j)}\right)=\mathcal{O}_{i}\left(P_{g,N}W_{(j)}\right)=\delta_{ij}$,
  4. (d)
    for any $R>0$, there are vector fields $Y_{(j)}^{a}\in W^{k,p}_{(2-n+\eta)^{+}}$ with $\eta=\max(\tau,-1)$ such that on $E_{R}$
    $W_{(j)}^{a}=G^{a}_{(j)}+Y_{(j)}^{a}.$ (3.11)
Proof.
Let $\chi(x)$ be a cutoff function that vanishes for $|x|>2$ and for $|x|<1/2$. We define $\chi_{R}(x):=\chi(x/R)$ and let $\widetilde{V}_{(j)}=r^{-1-n}R\chi_{R}e_{(j)}\in C_{c}^{\infty}$, where $R$ sufficiently large is chosen below. Using $U_{(i)}$ as introduced in the proof of Proposition 3.2, we define the quantity
$M_{ij}:=\int_{\mathbb{R}^{n}}\left<\widetilde{V}_{(j)},k_{(i)}\right>_{g}\;dV_{g}$ $=\int_{\mathbb{R}^{n}}r^{-1-n}R\chi_{R}\left<e_{(j)},e_{(i)}+U_{(i)}\right>_{g}\;dV_{g}$
$=c\delta_{ij}+o(1),$
where $c$ is a constant independent of $R$. Consequently $M_{ij}$ is invertible for $R$ sufficiently large: For some fixed such $R$, we take $V_{(i)}=\sum_{j}(M^{-1})_{ij}\widetilde{V}_{(j)}$ so that
$\int_{\mathbb{R}^{n}}\left<V_{(j)},k_{(i)}\right>_{g}\;dV_{g}=\delta_{ij}.$ (3.12)
We fix $\delta\in(2-n,0)$. Since $V_{(j)}\in C^{\infty}_{c}\subset W^{k-2,p}_{\delta-2}$, we can find a unique $W_{(j)}^{a}\in W^{k,p}_{\delta}$ with $(P_{g,N}W_{(j)})^{a}=V^{a}_{(j)}$, and as a consequence of uniqueness the vector field $W_{(j)}$ is independent of $\delta\in(2-n,0)$, establishing parts (a) and (b). Equation (3.12) precisely encodes that $\mathcal{O}_{i}(P_{g,N}W_{(j)})=\delta_{ij}$, so combining this with equation (3.7) establishes part (c).
To establish the decomposition (3.11) in part (d), we fix $R>0$ and observe that
$(\overline{P}W_{(j)})_{a}=\underbrace{(P_{g,N}W_{(j)})_{a}+((\overline{P}-P_{g,N})W)_{a}}_{=:F^{(j)}_{a}}$.
The first term defining $F^{(j)}_{a}$ is smooth and compactly supported, while the second term (using equation (3.2) and Lemma 2.4) lies in $W^{k-2,p}_{(\tau-n)^{+}}$. Hence $F^{(j)}_{a}\in W^{k-2,p}_{(\tau-n)^{+}}$, and Lemma A.9 implies that there exists $X_{(j)}^{a}\in W^{k,p}_{(2-n+\tau)^{+}}$ with $(\overline{P}X_{(j)})_{a}=F^{(j)}_{a}$ on $E_{R}$.
Since $\overline{P}(W_{(j)}-X_{(j)})=0$ on $E_{R}$, Lemma A.10 implies that there are constants $c_{ij}$ and vector fields $U_{(j)}^{a}\in W^{k,p}_{(1-n)^{+}}$ such that
$W_{(j)}^{a}-X_{(j)}^{a}=\sum_{i=1}^{n}c_{ij}G_{(i)}^{a}+U_{(j)}^{a}$
on $E_{R}$. Rearranging,
$W_{(j)}^{a}=\sum_{i=1}^{n}c_{ij}G_{(i)}^{a}+X_{(j)}^{a}+U_{(j)}^{a}.$ (3.13)
The decay rates of $X_{(j)}^{a}$ and $U_{(j)}^{a}$ imply that $Y_{(j)}^{a}:=X_{(j)}^{a}+U_{(j)}^{a}\in W^{k,p}_{(2-n+\eta)^{+}}$ with $\eta=\max(-1,\tau)$, so part (d) is proved up to determining the values of the constants $c_{ij}$.
To find $c_{ij}$, we plug (3.13) into the result of part (c), yielding
$\delta_{ij}=\mathcal{P}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W_{(j)}\right)=\lim_{R\to\infty}\int_{\partial B_{R}}\frac{1}{2N}\left[\sum_{k=1}^{n}c_{kj}(\operatorname{{\mathbb{L}}}G_{(k)})_{ab}+(\operatorname{{\mathbb{L}}}Y_{(j)})_{ab}\right]n^{b}e^{a}_{(i)}\;d\overline{A}.$
Since $\operatorname{{\mathbb{L}}}Y_{(j)}$ and $(\operatorname{{\mathbb{L}}}-\overline{\operatorname{{\mathbb{L}}}})G_{(k)}$ are $o(r^{1-n-\epsilon})$ for some $\epsilon>0$, and since $N-1=o(r^{\tau})$, we deduce that
$\delta_{ij}=\sum_{k=1}^{n}\frac{c_{kj}}{2}\left[\lim_{R\to\infty}\int_{\partial B_{R}}(\overline{\operatorname{{\mathbb{L}}}}G_{(k)})_{ab}n^{b}e^{a}_{(i)}\;d\overline{A}\right].$ (3.14)
Recalling equation (3.10), a computation shows that
$(\overline{\operatorname{{\mathbb{L}}}}G_{(k)})_{ab}=\frac{4n(n-2)C_{n}}{r^{n-1}}\left(\delta_{ak}n_{b}+\delta_{bk}n_{a}-\delta_{ab}n_{k}+(n-2)n_{a}n_{b}n_{k}\right),$ (3.15)
and contracting with $n^{b}$ yields
$(\overline{\operatorname{{\mathbb{L}}}}G_{(k)})_{ab}n^{b}=\frac{4n(n-2)C_{n}}{r^{n-1}}(\delta_{ak}+(n-2)n_{a}n_{k}).$
Since $\int_{\partial B_{R}}n_{a}n_{k}\;d\overline{A}=\frac{R^{n-1}}{n}|S^{n-1}|\delta_{ak}$, it follows that
$\int_{\partial B_{R}}(\overline{\operatorname{{\mathbb{L}}}}G_{(k)})_{ab}n^{b}e^{a}_{(i)}\;d\overline{A}=4n(n-2)|S^{n-1}|C_{n}\left(1+\frac{n-2}{n}\right)\delta_{ak}e^{a}_{(i)}=2\delta_{ik}.$
Recalling (3.14), we find that $c_{ij}=\delta_{ij}$, which completes the proof of part (d).
Armed with the momentum carrier vector fields $W_{(j)}^{a}$, we obtain the following precise description of the solution of equation (3.1) if the right-hand side is rapidly decaying:
Proposition 3.5.
Suppose that $g_{ab}$ and $N$ satisfy Assumption 2.3 with parameters $k$, $p$, and $\tau$. Assume $1-n<\delta<2-n$, and let $Z_{b}\in W^{k-2,p}_{\delta-2}$. Recall that
$\mathcal{O}_{i}(Z)=\int Z_{b}k^{b}_{(i)}\;dV_{g}$
and define
$U_{b}:=Z_{b}-\sum_{i=1}^{n}\mathcal{O}_{i}(Z)(P_{g,N}W_{(i)})_{b}.$ (3.16)
Then there exists a unique $V^{a}\in W^{k,p}_{\delta}$ solving
$(P_{g,N}V)_{a}=U_{b}.$ (3.17)
Hence
$W^{a}=\underbrace{\sum_{i=1}^{n}\mathcal{O}_{i}(Z)W_{(i)}^{a}}_{O(r^{2-n})}+\underbrace{V^{a}}_{o(r^{2-n}))}$ (3.18)
is the unique solution in $W^{k,p}_{(2-n)^{+}}$ of equation (3.1). Moreover, the ADM momentum of $1/(2N)(\operatorname{{\mathbb{L}}}W)_{ab}$ satisfies
$\mathcal{P}_{i}\left(\frac{1}{2N}(\operatorname{{\mathbb{L}}}W)\right)=\mathcal{O}_{i}(Z),\quad i=1,\ldots,n.$
Proof.
Since each $(P_{g,N}W_{(j)})_{a}$ is smooth and compactly supported, $U_{b}$ defined by equation (3.16) has the same regularity and decay properties as $Z_{b}$. That is, $U_{b}\in W^{k-2,p}_{\delta-2}$. Moreover, since $\mathcal{O}_{i}(P_{g,N}W_{(j)})=\delta_{ij}$ we find that
$\mathcal{R}_{i}(U)=\mathcal{O}_{i}(Z)-\sum_{j=1}^{n}\mathcal{R}_{j}(Z)\mathcal{R}_{i}(P_{g,N}W_{(j)})=0.$
Since the obstruction coefficients vanish, Proposition 3.1(f) provides a unique solution $V^{a}\in W^{k,p}_{\delta}$ of (3.17). Hence $W^{a}$ defined by (3.18) is the unique solution in $W^{k,p}_{(2-n)^{+}}$ of equation (3.1) provided by Proposition 3.1. Finally, using Lemma 3.3 we compute the ADM momentum
$\mathcal{P}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W\right)=\mathcal{R}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W\right)+\mathcal{G}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W\right)=\mathcal{O}_{i}(P_{g,N}W)=\mathcal{O}_{i}(Z).$
Combining Proposition 3.5 and Proposition 3.4(d) we find that if $Z_{a}$ has fast decay, $o(r^{-n})$, then the solution of equation (3.1) has a leading order $O(r^{2-n})$ term that is a linear combination of the columns of the Green’s function of the Euclidean vector Laplacian, with a correction decaying at a faster rate depending on both $\tau$ and $\delta$. We now investigate the borderline case such that $Z_{a}$ decays at the threshold rate $r^{-n}$. The situation is more delicate because the vector Laplacian is no longer Fredholm as a map on the full weighted space $W^{k,p}_{2-n}$. To compensate for this, we confine our attention to the case such that $Z_{a}$ admits the expansion $Z_{a}=r^{-n}F_{a}+H_{a}$ on some exterior region $E_{R}$, where $F_{a}$ is homogeneous of degree zero and where $H_{a}$ lies in a weighted space with decay faster than $r^{-n}$. In this scenario, we seek a solution of the form $W^{a}=r^{2-n}U^{a}+V^{a}$, where $U^{a}$ is homogeneous of degree zero and where $V^{a}$ decays faster than $r^{2-n}$. The following result concerning the Euclidean vector Laplacian is the tool needed for the solvability for the leading order homogeneous terms.
Proposition 3.6.
Take $k\geq 2$ and $1<p<\infty$. Suppose that the covector field $Z_{a}\in W^{k-2,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ has the form $Z_{a}=r^{-n}F_{a}$, where $F_{a}\in W^{k-2,p}(S^{n-1})$ is extended homogeneously with degree zero. Then there exists a vector field $W^{a}\in W^{k,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ of the form $W^{a}=r^{2-n}U^{a}$, where $U^{a}\in W^{k,p}(S^{n-1})$ is extended homogeneously with degree zero, solving $(\overline{P}W)_{a}=Z_{a}$ if and only if
$\int_{S^{n-1}}F(e_{(i)})\;d\overline{A}=0\quad\text{for}\quad i=1,\ldots,n.$ (3.19)
The solution $W^{a}$ is unique up to a linear combination of the vector fields $G^{a}_{(j)}$.
Proof.
Consider a smooth vector field $W^{a}$ on $\mathbb{R}^{n}\setminus\{0\}$ of the form $r^{\alpha}(V^{a}+vn^{a})$ where $\alpha\in\mathbb{R}$, and where $V^{a}$ is a vector field in $W^{k,p}(S^{n-1})$ and tangential to $S^{n-1}$ and $v$ is a function in $W^{k,p}(S^{n-1})$, both extended to $\mathbb{R}^{n}\setminus\{0\}$ by zero-homogeneity. An extended computation shows that $(\overline{P}W)^{a}=\frac{1}{2}r^{\alpha-2}(F^{a}+fn^{a})$ where $F^{a}$ and $f$ are a tangential vector field and function respectively in $W^{k-2,p}(S^{n-1})$ with
$\begin{bmatrix}F^{a}\\ f\end{bmatrix}=\begin{bmatrix}\not{\Delta}+\left(1-\frac{2}{n}\right)\not{\nabla}\not{\operatorname{\mathrm{div}}}+a_{\alpha}&b_{\alpha}\not{\nabla}\\ c_{\alpha}\not{\operatorname{\mathrm{div}}}&\not{\Delta}+d_{\alpha}\end{bmatrix}\begin{bmatrix}V^{a}\\ v\end{bmatrix}$.
Here, the slashed operators are intrinsic to the round $S^{n-1}$ and the coefficients are
$a_{\alpha}$ $=\alpha(\alpha+n-2)-1$
$b_{\alpha}$ $=n+(\alpha-1)\left(1-\frac{2}{n}\right)$
$c_{\alpha}$ $=\alpha-3-\frac{2}{n}(\alpha-1)$
$d_{\alpha}$ $=2\frac{(\alpha-1)(n-1)}{n}(\alpha+n-1)$
The operators $L_{\alpha}$ sending $(V^{a},v)\mapsto(F^{a},f)$ as above are elliptic and Fredholm as maps $W^{k,p}\to W^{k-2,p}$ on the bundle over $S^{n-1}$ with sections being pairs of tangential vector fields and functions mapping $S^{n-1}\to\mathbb{R}$. These operators are homotopic via Fredholm maps to $\mathrm{diag}(\not{\Delta},\not{\Delta})$, which is formally self adjoint and hence has index zero. The maps $L_{\alpha}$ then also have index zero. They are not self adjoint in general, but since
$a_{0}=a_{2-n};\quad b_{0}=-c_{2-n};\quad c_{0}=-b_{2-n};\quad d_{0}=d_{2-n},$
and since $(\not{\operatorname{\mathrm{div}}})^{*}=$-$\not{\nabla}$, it follows that the formal adjoint of $L_{2-n}$ is $L_{0}$. Hence, given a choice of $(F^{a},f)\in W^{k-2,p}(S^{n-1})$ the equation $L_{2-n}(V^{a},v)=(F^{a},f)$ is solvable for $(V^{a},v)\in W^{k,p}(S^{n-1})$ if and only if $(F^{a},f)$ is $L^{2}$ orthogonal to the kernel of $L_{0}$.
To identify this kernel, it is helpful to return to Cartesian coordinates, wherein the action of $\overline{P}$ takes the form
$2(\overline{P}X)_{a}=\overline{\Delta}(X_{a})+\left(1-\frac{2}{n}\right)\overline{\nabla}_{a}(\overline{\nabla}_{b}X^{b}),$
where $\overline{\Delta}$ applies component-wise on the right hand side. We are then seeking vector fields $X^{a}$ that are homogeneous of degree 0 and are contained in the kernel of $\overline{P}$. We claim that such a vector field must be divergence-free, satisfying $\overline{\nabla}_{b}X^{b}=0$. Assuming this is true, then
$0=2(\overline{P}X)_{a}=\overline{\Delta}(X_{a}).$
Hence the components of $X^{a}$ are harmonic and homogeneous of degree 0, and it follows that they are constants. Conversely, constant vector fields are clearly contained in the kernel, so we conclude that the kernel of $L_{0}$ precisely consists of the constant vector fields on $\mathbb{R}^{n}$, restricted to the sphere and decomposed into tangential and normal components. The solvabilty conditions (3.19) then follow.
To establish the claim that $\overline{\nabla}_{b}X^{b}=0$, we observe that
$0=\overline{\nabla}^{a}(\overline{P}X)_{a}=\left(1-\frac{1}{n}\right)\overline{\Delta}(\overline{\nabla}_{b}X^{b}).$
Since $1-(1/n)\neq 0$, we conclude that $\overline{\nabla}_{b}X^{b}$ is harmonic and homogeneous of degree 1-1. If $n>3$ there are no such nonzero functions since 2n<1<02-n<-1<0, which establishes the claim in this case. Now suppose that $n=3$, in which case $\overline{\nabla}_{b}X^{b}=c/r$ for some constant $c$. We need to show $c=0$.
Since $(\overline{P}X)_{a}=0$ and since $\overline{\nabla}_{b}X^{b}=c/r$ we have
$\overline{\Delta}(X_{a})=-\left(1-\frac{2}{3}\right)\overline{\nabla}_{a}(c/r)=\frac{c}{3r^{2}}n_{a}.$
Using the fact that the first spherical harmonics on $S^{2}$ (i.e., the restrictions of the components of $n^{a}$ to the sphere) all have eigenvalue 2-2 under the Laplacian, we find that $X^{a}=C^{a}-\frac{c}{6}n^{a}$ for some constant vector field $C^{a}$. Taking a divergence, we compute that $\overline{\nabla}_{b}X^{b}=-c/(3r)$. Recalling that we began with $\overline{\nabla}_{b}X^{b}=c/r$, we conclude that $c=0$, as required.
To finish the proof, it remains to identify the kernel of $L_{2-n}$. The $n$ vector fields $G^{a}_{(j)}$ from the Greens function of $\overline{P}$ are homogeneous of degree 2n2-n, and each provides an element of this kernel. Since $L_{2-n}$ has index zero and a cokernel of dimension $n$, its kernel has dimension $n$, so these provide a complete basis.
With this result for the Euclidean operator $\overline{P}$ in hand, we can now solve the broader problem for $P_{g,N}$, supposing that our leading order source term $F_{a}$ satisfies the solvability condition:
Proposition 3.7.
Suppose that $g_{ab}$ and $N$ satisfy Assumption 2.3 with parameters $k$, $p$, and $\tau$. Suppose that the covector field $Z_{a}\in W^{k-2,p}_{(-n)^{+}}$ admits the decomposition
$Z_{a}=r^{-n}F_{a}+H_{a}$
on $E_{R}$ for some $R>0$, where $H_{a}\in W^{k-2,p}_{\delta-2}$ for some $1-n<\delta<2-n$, and where $F_{a}\in W^{k-2,p}(S^{n-1})$ is extended homogeneously with degree zero. If
$\int_{S^{n-1}}F(e_{(i)})\;d\overline{A}=0\quad\text{for}\quad i=1,\ldots,n,$ (3.20)
then the unique solution $W^{a}\in W^{k,p}_{(2-n)^{+}}$ of $(P_{g,N}W)_{a}=Z_{a}$ admits the decomposition
$W^{a}=r^{2-n}U^{a}+V^{a}$ (3.21)
on $E_{R}$, where $V^{a}\in W^{k,p}_{\delta}\cup W^{k,p}_{(2-n+\tau)^{+}}$ and $U_{a}\in W^{k,p}(S^{n-1})$ is extended homogeneously with degree zero.
Proof.
Proposition 3.6 ensures that we may find $\overline{W}^{a}\in W^{k,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ of the form $\overline{W}^{a}=r^{2-n}\overline{U}^{a}$, with $\overline{U}^{a}\in W^{k,p}(S^{n-1})$ extended homogeneously with degree zero, satisfying $(\overline{P}\,\overline{W})_{a}=r^{-n}F_{a}$. This determines $\overline{W}^{a}$ up to a linear combination of the $G^{a}_{(j)}$ vector fields– we fix here one particular choice.
Let $\chi$ be a cutoff function that equals 0 on $B_{R/2}$ and $1$ on $E_{R}$. Then $\overline{V}^{a}:=W^{a}-\chi\overline{W}^{a}\in W^{k,p}_{(2-n)^{+}}$ satisfies
$(P_{g,N}\overline{V})_{a}$ $=(P_{g,N}W)_{a}-(\overline{P}\,\chi\overline{W})_{a}-((P_{g,N}-\overline{P})\chi\overline{W})_{a}$
$=H_{a}-((P_{g,N}-\overline{P})\chi\overline{W})_{a}$
on $E_{R}$. Since $\chi\overline{W}^{a}\in W^{k,p}_{(2-n)^{+}}$, the latter term is in $W^{k-2,p}_{(\tau-n)^{+}}$, so that $(P_{g,N}\overline{V})_{a}\in W^{k-2,p}_{\delta-2}\cup W^{k-2,p}_{(\tau-n)^{+}}$. As a consequence of Propositions 3.5 and 3.4(d), $\overline{V}^{a}$ admits the decomposition
$\overline{V}^{a}=\sum_{j=1}^{n}\mathcal{O}_{j}(P_{g,N}\overline{V})G^{a}_{(j)}+V^{a}$
on $E_{R}$, where $V^{a}\in W^{k,p}_{\delta}\cup W^{k,p}_{(2-n+\tau)^{+}}$. Since we have on $E_{R}$ that
$W^{a}=\overline{W}^{a}+\overline{V}^{a}=r^{2-n}\overline{U}^{a}+\sum_{j=1}^{n}\mathcal{O}_{j}(P_{g,N}\overline{V})G^{a}_{(j)}+V^{a},$
this establishes (3.21) with $U^{a}=\overline{U}^{a}+\sum_{j=1}^{n}\mathcal{O}_{j}(P_{g,N}\overline{V})r^{n-2}G^{a}_{(j)}$.
We now show that the obstruction criteria (3.20) of the previous proposition are automatically satisfied for our primary use case where the source term is a divergence.
Lemma 3.8.
Suppose $S_{ab}\in W^{k-1,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ is homogeneous of degree 1n1-n. Then
$\overline{\nabla}^{a}S_{ab}=r^{-n}F_{b}$ (3.22)
for some $F_{b}\in W^{k-2,p}(S^{n-1})$ extended homogeneously with degree zero and satisfying
$\int_{S^{n-1}}F(e_{(i)})\;d\overline{A}=0\quad\text{for}\quad i=1,\ldots,n.$ (3.23)
Proof.
The form for $\overline{\nabla}^{a}S_{ab}$ given by equation (3.22) is a fairly direct consequence of homogeneity, so it remains to establish the vanishing of the obstruction coefficients in equation (3.23) .
Consider $r_{1},r_{2}\in\mathbb{R}$ with $0<r_{1}<r_{2}$ and let $A$ be the annulus bounded by $r_{1}<r<r_{2}$. Then
$\int_{A}\overline{\nabla}^{a}(S_{ab})e^{b}_{(i)}\;dV_{\overline{g}}=\int_{r_{1}}^{r_{2}}r^{-1}\int_{S^{n-1}}F_{a}e^{a}_{(i)}\;d\overline{A}\;dr=\log(r_{2}/r_{1})\int_{S^{n-1}}F_{a}e^{a}_{(i)}\;d\overline{A}.$
On the other hand, the divergence theorem implies that
$\int_{A}\overline{\nabla}^{a}(S_{ab})e^{b}_{(i)}\;dV_{\overline{g}}$ $=\int_{A}\overline{\nabla}^{a}(S_{ab}e^{b}_{(i)})\;dV_{\overline{g}}=\int_{\partial B_{r_{2}}}S_{ab}e^{b}_{(i)}n^{a}\;dA_{\overline{g}}-\int_{\partial B_{r_{1}}}S_{ab}e^{b}_{(i)}n^{a}\;dA_{\overline{g}}$
$=\int_{S^{n-1}}B_{ab}e^{b}_{(i)}n^{a}\;d\overline{A}-\int_{S^{n-1}}B_{ab}e^{b}_{(i)}n^{a}\;d\overline{A}=0.$
Equating the final expressions from the two computations above, we find that
$\log(r_{2}/r_{1})\int_{S^{n-1}}F_{a}e^{a}_{(i)}\;d\overline{A}=0,$
which establishes the result.
Compiling the results in this section, we have the following theorem summarizing the prescription of the asymptotics of the unphysical second fundamental form $K_{ab}$ in the conformal method:
Theorem 3.9.
Suppose that $g_{ab}$ and $N$ satisfy Assumption 2.3 with parameters $k$, $p$, and $\tau$. We take $A_{ab}\in W^{k-1,p}_{\delta-1}$ for some $\delta<0$ to be symmetric and trace-free, and we define $Z_{b}=\nabla^{a}\left(\frac{1}{2N}A_{ab}\right)$. Given $W^{a}$ solving $(P_{g,N}W)_{a}=Z_{a}$, we set $K_{ab}=1/(2N)(A_{ab}-(\operatorname{{\mathbb{L}}}W)_{ab})$ and $S_{ab}=1/(2N)A_{ab}$.
  1. (a)
    If $2-n<\delta<0$, then there exists a unique $W^{a}\in W^{k,p}_{\delta}$ solving $(P_{g,N}W)_{a}=Z_{a}$.
  2. (b)
    If $1-n<\delta<2-n$, then the unique $W^{a}\in W^{k,p}_{(2-n)^{+}}$ solving $(P_{g,N}W)_{a}=Z_{a}$ may be expanded in terms of the momentum carriers,
    $W^{a}=\sum_{i=1}^{n}\mathcal{O}_{i}(Z)W^{a}_{(i)}+V^{a},$
    with the remainder $V^{a}\in W^{k,p}_{\delta}$. Moreover, $\mathcal{P}_{i}(K)=\mathcal{G}_{i}(S)=\mathcal{P}_{i}(S)-\mathcal{R}_{i}(S)=-\mathcal{O}_{i}(Z)$.
  3. (c)
    In the borderline case such that $A_{ab}\in W^{k-1,p}_{(1-n)^{+}}$ has the form
    $A_{ab}=r^{1-n}B_{ab}+C_{ab}$ (3.24)
    on $E_{R}$ for some $R>0$, where $B_{ab}\in W^{k-1,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ is homogeneous of degree zero and $C_{ab}\in W^{k-1,p}_{\delta-1}$ for some $1-n<\delta<2-n$ (with $B_{ab}$ and $C_{ab}$ both traceless and symmetric), the unique $W^{a}\in W^{k,p}_{(2-n)^{+}}$ solving $(P_{g,N}W)_{a}=Z_{a}$ is given by
    $W^{a}=r^{2-n}U^{a}+V^{a}$ (3.25)
    on $E_{R}$, where $U^{a}\in W^{k,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ is homogeneous of degree zero and $V^{a}\in W^{k,p}_{\delta}\cup W^{k,p}_{(2-n+\tau)^{+}}$. Moreover, $\mathcal{P}_{i}(K)=\mathcal{G}_{i}(S)=\mathcal{P}_{i}(S)-\mathcal{R}_{i}(S)$.
Proof.
Part (a) follows immediately from Proposition 3.1(e).
Recalling that $\mathcal{G}_{i}(S)=\mathcal{P}_{i}(S)-\mathcal{R}_{i}(S)$ holds definitionally, the momentum identifications $\mathcal{P}_{i}(K)=\mathcal{G}_{i}(S)=\mathcal{P}_{i}(S)-\mathcal{R}_{i}(S)$ in both parts (b) and (3.25) follow from expanding $\mathcal{P}_{i}(K)$ into $\mathcal{G}_{i}(K)$ and $\mathcal{R}_{i}(K)$ and using that $\mathcal{R}_{i}(K)=0$ (which holds because $\nabla^{a}K_{ab}=0$) and that $\mathcal{G}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W\right)=0$ (from Lemma 3.3). Explicitly:
$\mathcal{P}_{i}(K)=\mathcal{G}_{i}(K)+\mathcal{R}_{i}(K)=\mathcal{G}_{i}(S)-\mathcal{G}_{i}\left(\frac{1}{2N}\operatorname{{\mathbb{L}}}W\right)=\mathcal{G}_{i}(S),$
The decomposition of $W^{a}$ in part (b) follows immediately from Proposition 3.5. The further momentum identification $\mathcal{P}_{i}(K)=-\mathcal{O}_{i}(Z)$ in (b) is due to the fact that the fast decay of $S_{ab}$ results in $\mathcal{P}_{i}(S)=0$, and from the definitional relation $\mathcal{R}_{i}\left(S\right)=\mathcal{O}_{i}(Z)$.
The decomposition in part (3.25) follows from Proposition 3.7 since Lemma 3.8 ensures that $\overline{\nabla}^{a}(\frac{1}{2N}r^{1-n}B_{ab})=r^{-n}F_{b}$ for some $F_{b}\in W^{k-2,p}(S^{n-1})$ satisfying (3.23) and which is extended homogeneously with degree zero, so that on $E_{R}$ we may write
$Z_{b}=\nabla^{a}\left(\frac{1}{2N}A_{ab}\right)=\overline{\nabla}^{a}\left(\frac{1}{2N}r^{1-n}B_{ab}\right)+(\nabla^{a}-\overline{\nabla}^{a})\left(\frac{1}{2N}r^{1-n}B_{ab}\right)+\nabla^{a}\left(\frac{1}{2N}C_{ab}\right)=r^{-n}F_{b}+H_{b},$
where we have (letting $\chi$ be a cutoff function that equals 0 on $B_{R/2}$ and $1$ on $E_{R}$)
$H_{b}:=\chi(\nabla^{a}-\overline{\nabla}^{a})\left(\frac{1}{2N}r^{1-n}B_{ab}\right)+\nabla^{a}\left(\frac{1}{2N}C_{ab}\right)\in W^{k-2,p}_{\delta-2}\cup W^{k-2,p}_{(\tau-n)^{+}}.$
This summary indicates the precise extent of control one has over the structure of the unphysical second fundamental form $K_{ab}$ with regards to its ADM momentum and decay properties. In particular, if one has sufficient decay to make sense of each of $\mathcal{G}_{i}(S)$ and $\mathcal{R}_{i}(S)$, the ADM momentum components are $\mathcal{P}_{i}(K)=\mathcal{G}_{i}(S)=\mathcal{P}_{i}(S)-\mathcal{R}_{i}(S)$. Given conformal method seed data $g_{ab}$ and $N$ satisfying Assumption 2.3 and $A_{ab}$ with at least $O(r^{1-n})$ decay, $\mathcal{G}_{i}(A_{ab}/2N)$ therefore directly measures the ADM momentum of the eventual initial data set, and it follows that this momentum can be readily prescribed a priori. Moreover, the quantity $\mathcal{R}_{i}(A_{ab}/2N)$ specifies the momentum which must be removed to satisfy the momentum constraint. These results motivate our use of the monikers “gravitational momentum” for $\mathcal{G}_{i}$ and “residual momentum” for $\mathcal{R}_{i}$.
In all cases, one can precisely control the decay rate of $K_{ab}$ (beyond any $r^{1-n}$ terms that arise) by appropriately choosing the decay rates of one’s seed data. To illustrate, we now discuss the precise decay structure of $W^{a}$ and $K_{ab}$ which we obtain in solving (3.1) in the fast-decay setting of $A_{ab}\in W^{k-1,p}_{\delta-1}$ with $1-n<\delta<2-n$, part (b) above. Abbreviating $p_{i}:=\mathcal{G}_{i}(S)=-\mathcal{O}_{i}(Z)$, we find that $W^{a}$ contains an $O(r^{2-n})$ piece given by $\sum_{j}-p_{j}W_{(j)}^{a}$, and the remainder $V^{a}$ beyond this has the desired $o(r^{\delta})$ decay inherited from $A_{ab}\in W^{k-1,p}_{\delta-1}$. In light of Proposition 3.4(d), however, one can be more explicit in describing the asymptotic form of $W^{a}$: For every $\epsilon>0$ and any $\ell\in\mathbb{N}$ such that $k-1>\ell+n/p$, one has
$W^{a}=\sum_{j=1}^{n}-p_{j}G_{(j)}^{a}+o_{\ell+1}(r^{\max(\delta,\,2-n+\tau+\epsilon)}),$
and the corresponding asymptotics of $K_{ab}$ are the following (using (3.15)):
$K_{ab}$ $=\frac{1}{2}\sum_{j=1}^{n}p_{j}\cdot(\overline{\operatorname{{\mathbb{L}}}}G_{(j)})_{ab}+o_{\ell}(r^{\max(\delta-1,\,1-n+\tau+\epsilon)})$
$=\frac{2n(n-2)C_{n}}{r^{n-1}}\left[p_{a}n_{b}+p_{b}n_{a}+((n-2)n_{a}n_{b}-\delta_{ab})p_{j}n^{j}\right]+o_{\ell}(r^{\max(\delta-1,\,1-n+\tau+\epsilon)}).$ (3.26)
In particular, for $n=3$ one has
$K_{ab}=\frac{3}{16\pi r^{2}}\left[p_{a}n_{b}+p_{b}n_{a}+(n_{a}n_{b}-\delta_{ab})p_{j}n^{j}\right]+o_{\ell}(r^{\max(\delta-1,\,-2+\tau+\epsilon)}).$
In the case of fast decay, we may therefore anticipate the complete structure of the leading $O(r^{1-n})$ term, and we may dictate the decay rate of $K_{ab}$, and hence of $\widetilde{K}_{ab}$, beyond this term through the decay of the provided seed data $g_{ab}$ and $N$ (specifying $\tau$) and $A_{ab}$ (specifying $\delta$). One obtains as many derivatives of control in the error term as desired by choosing the seed data to allow $k$ sufficiently large. Although it is not typically of physical interest, we remark that one could obtain similar results down to $n<\delta<1-n$, or for an arbitrarily negative value of $\delta$, by expanding Proposition 3.2 to characterize the kernel of $P_{g,N}$ on higher-weighted spaces and expanding Proposition 3.4 to introduce the faster-decaying carrier vector fields dual to these additional kernel elements.
While we can anticipate the complete structure of the leading $O(r^{1-n})$ term of $K_{ab}$ in the case of fast decay discussed above, there is more freedom if $A_{ab}$ is itself given with the threshold $O(r^{1-n})$ decay, as in part (3.25) of Theorem 3.9. A question then remains regarding what kinds of leading terms $r^{1-n}D_{ab}$ ($D_{ab}$ is analogous to the quantity in brackets in equation (3.26)) are possible for $K_{ab}$ in this case. The answer is that one can choose any Euclidean transverse traceless tensor of this form: the possible homogeneous of degree zero tensors $D_{ab}\in W^{k-1,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ are precisely those which are symmetric and traceless and satisfy $\overline{\nabla}^{a}(r^{1-n}D_{ab})=0$. This form is necessary because isolating the leading order $r^{1-n}D_{ab}$ term in $K_{ab}=\frac{1}{2N}(A_{ab}-(\operatorname{{\mathbb{L}}}W)_{ab})$, with $A_{ab}$ and $W^{a}$ as in (3.24) and (3.25) on some $E_{R}$, yields that symmetry and tracelessness are immediate from these properties of $K_{ab}$ overall, and the leading order term in $\nabla^{a}K_{ab}=0$ reads $\overline{\nabla}^{a}(r^{1-n}D_{ab})=0$.
Conversely, beginning with a given Euclidean transverse traceless tensor $r^{1-n}D_{ab}$, if one sets
$A_{ab}=2N\chi r^{1-n}D_{ab}\in W^{k-1,p}_{(1-n)^{+}}$
with $\chi$ a cutoff function equal to 0 on $B_{R/2}$ and $1$ on some $E_{R}$, then on $E_{R}$ one has
$Z_{b}:=\nabla^{a}\left(\frac{1}{2N}A_{ab}\right)=\nabla^{a}\left(r^{1-n}D_{ab}\right)=(\nabla^{a}-\overline{\nabla}^{a})\left(r^{1-n}D_{ab}\right)=o(r^{\tau-n}),.$
This result implies that $Z_{b}\in W^{k-2,p}_{(-n)^{-}}$, and consequently Proposition 3.5 yields the decomposition (3.18) of the solution $W^{a}\in W^{k,p}_{(2-n)^{+}}$. The leading order $O(r^{1-n})$ terms in the expansion of $K_{ab}=\frac{1}{2N}(A_{ab}-(\operatorname{{\mathbb{L}}}W)_{ab})$ are then $r^{1-n}D_{ab}$ and the leading term in equation (3.26), with $p_{j}=-\mathcal{R}_{j}(r^{1-n}D_{ab})$. This extraneous term can be removed by perturbing $A_{ab}$ at lower order (specifically, by subtracting some $B_{ab}\in W^{k-1,p}_{(1-n)^{-}}$ satisfying $\mathcal{R}_{j}(\frac{1}{2N}B_{ab})=\mathcal{R}_{j}(r^{1-n}D_{ab})$), resulting in $K_{ab}$ having the leading order term $r^{1-n}D_{ab}$, as desired.
We remark that the problem of characterizing and constructing Euclidean transverse traceless tensors is well-studied in the literature beig1997tt ; tafel2018all , so numerous such objects can be readily constructed. The procedure outlined in part (3.25) of Theorem 3.9 and the above discussion then allows one to build any example into a working $K_{ab}$ as the leading order $O(r^{1-n})$ term if desired.
4 The Hamiltonian Constraint
Assured that we may construct a continuous $K_{ab}$ with any requisite asymptotics, we suppose now that we have done this and turn to the vacuum Hamiltonian constraint, which in the conformal method takes the form of the Lichnerowicz equation (2.5),
$(q_{n}+2)\Delta_{g}\varphi+R(g)\varphi=\varphi^{-q_{n}-1}|K|_{g}^{2}$ (4.1)
(we recall $q_{n}:=\frac{2n}{n-2}$), to be solved for the conformal factor $\varphi>0$ satisfying $\lim_{|x|\to\infty}\varphi=1$. Despite the nonlinearity of this equation, the problem of existence and uniqueness of solutions has been thoroughly resolved in the literature bruhat-isenberg ; maxwell_solutions_2005 . One first observes that, in addition to Assumption 2.3, it is necessary to require that $g$ is Yamabe positive, which means that the conformal invariant
[…]
is strictly positive. To see that this is a necessary condition, we suppose that there exists a positive solution $\varphi$ to equation (4.1) with $\varphi-1\in W^{k,p}_{\delta}$ for some $\delta<0$. Under the conformal change specified by $\varphi$, one has $Y_{\tilde{g}}>0$ because $R(\tilde{g})=|K|_{\tilde{g}}^{2}$ is positive and continuous, and because Sobolev embedding ensures that
$\|f\|_{L^{q_{n}}}^{2}\lesssim\|\nabla f\|_{L^{2}}^{2}+\|R(\tilde{g})^{1/2}f\|_{L^{2}}^{2}.$
The notation $\lesssim$” here indicates that one has a standard inequality, with $\leq$”, up to an overall positive multiplicative constant independent of the quantities of interest (in this case, of $f\in C_{c}^{\infty}$). Conformal invariance now yields $Y_{g}>0$.
It turns out that $Y_{g}>0$ is also a sufficient condition, as indicated by the following existence theorem for solutions of the Lichnerowicz equation for asymptotically Euclidean seed data. This theorem is implied by Theorem 4.4 of maxwell_solutions_2005 ; for uniqueness, see the discussion in Section VIII of bruhat-isenberg .
Theorem 4.1.
Suppose that $g_{ab}$ satisfies Assumption 2.3 with parameters $k$, $p$, and $\tau$ with $2-n<\tau<0$, and that $K_{ab}\in W^{k-1,p}_{\tau-1}$. If $Y_{g}>0$, then there exists a unique positive conformal factor $\varphi$ solving the Lichnerowicz equation (4.1) with $\varphi-1\in W^{k,p}_{\tau}$.
It is useful to also state precisely the implication of this result in the case of faster decay:
Corollary 4.2.
Suppose that $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{(2-n)^{+}}$ with $k>1+n/p$, and that $K_{ab}\in W^{k-1,p}_{(1-n)^{+}}$. If $Y_{g}>0$, then there exists a unique positive conformal factor $\varphi$ solving the Lichnerowicz equation (4.1) with $\varphi-1\in W^{k,p}_{(2-n)^{+}}$.
As in Section 3, understanding the asymptotics of these solutions requires a finer analysis of the next decay range $1-n<\delta<2-n$. Since we are now dealing with the simpler scalar Laplacian rather than the vector Laplacian, much of the requisite analysis is a simplified version of that developed in Section 3 and is well understood in the literature. To make the relation between the analyses of these two Laplacians explicit, we present this theory in a manner closely analogous to the discussion in Section 3. To most naturally incorporate the boundary condition $\varphi\to 1$, we write $\varphi=1+u$ and recast (4.1) as follows:
$(q_{n}+2)\Delta_{g}u+(1+u)R(g)=(1+u)^{-q_{n}-1}|K|_{g}^{2}$ (4.3)
This is the equation we analyze. We begin our analysis by noting the close connection between $\Delta_{g}$ and the Euclidean Laplacian $\overline{\Delta}$ induced by the following equation:
$\Delta_{g}u=\overline{\Delta}u+\sum_{|\alpha|=1}^{2}h^{\alpha}\partial_{\alpha}u,$ (4.4)
with each $h^{\alpha}\in W^{k-2+|\alpha|,p}_{\tau-2+|\alpha|}$. Since the following statement is well known, and since the details are much the same as those of Proposition 3.1, we omit the proof.
Proposition 4.3.
Suppose that $g_{ab}$ satisfies Assumption 2.3 with parameters $k$, $p$, and $\tau$, and that $V\in W^{k-2,p}_{\tau-2}$. If $\delta$ is non-exceptional, then $L=-\Delta_{g}+V:W^{k,p}_{\delta}\to W^{k-2,p}_{\delta-2}$ is continuous and Fredholm, with its index equal to that of the Euclidean Laplacian acting between the same spaces. If $V\geq 0$, then $L$ is an isomorphism for $2-n<\delta<0$.
The following proposition and corollary are a variation of Proposition 1 in bieri_brill_2025 in that they extract the mass term from $\varphi$ to control the remaining decay. Following the thread of Section 3, we reformulate this as relying on the existence of a “mass carrier” function– compare this to Proposition 3.4 (the constant function $1$ now plays the role analogous to that of the kernel vector fields $k_{(j)}^{a}$).
Proposition 4.4.
Suppose that $g_{ab}$ satisfies Assumption 2.3 with parameters $k$, $p$, and $\tau$. There exists a function $u_{0}\in W^{k,p}_{(2-n)^{+}}$ satisfying:
  1. (a)
    $\Delta_{g}u_{0}$ is smooth, nonnegative, and compactly supported,
  2. (b)
    $\int_{\mathbb{R}^{n}}(-\Delta_{g}u_{0})\;dV_{g}=1$,
  3. (c)
    for each $R>0$, there is a function $v\in W^{k,p}_{(2-n+\eta)^{+}}$ with $\eta=\max(\tau,-1)$ such that
    $u_{0}=\frac{r^{2-n}}{(n-2)|S^{n-1}|}+v$ (4.5)
    on the exterior region $E_{R}$.
Proof.
Let $\mu$ be a smooth, non-negative, compactly supported function with $\int_{\mathbb{R}^{n}}\mu\;dV_{g}=1$. Hence $\mu\in W^{k-2,p}_{\delta}$ for every $\delta$, so Proposition 4.3 implies that we can find $u_{0}\in W^{k,p}_{(2-n)^{+}}$ with $\Delta_{g}u_{0}=\mu$. We observe that
$\overline{\Delta}u_{0}=\underbrace{\Delta_{g}u_{0}+(\overline{\Delta}-\Delta_{g})u_{0}}_{=:F},$
where the first term contained in $F$ is smooth and compactly supported, and the second term (and hence $F$ itself) belongs to $W^{k-2,p}_{(\tau-n)^{+}}$ as a consequence of equation (4.4) and Lemma 2.4. Lemma 1 of bieri_brill_2025 implies that, for each $R>0$, there exists a function $w\in W^{k,p}_{(2-n+\tau)^{+}}$ with $\overline{\Delta}w=F$ on $E_{R}$. Thus $h:=u_{0}-w$ satisfies $\overline{\Delta}h=0$ on $E_{R}$. Since $h$ decays as $o(1)$, the classical multipole expansion for harmonic functions shows that
$h=\alpha r^{2-n}+\delta h$
on $E_{R}$, where $\alpha$ is a constant and $\delta h$ is a harmonic function of order 1n1-n on $\mathbb{R}^{n}\setminus\{0\}$.
Now choose $\chi$ to be a cutoff function that equals 0 on $B_{R/2}$ and 1 on $E_{R}$. We set
$v=\chi w+\chi\delta h,$
and we observe that the decay rates of the two terms imply that $v\in W^{k,p}_{(2-n+\eta)^{+}}$ with $\eta=\max(\tau,-1)$. Since $u_{0}=\alpha r^{2-n}+v$ on $E_{R}$, it remains only to determine the value of $\alpha$. We integrate by parts and invoke the fast decay of $v$ to find that
$1=\int_{\mathbb{R}^{n}}(-\Delta_{g}u_{0})\;dV_{g}=\lim_{R^{\prime}\to\infty}\int_{\partial B_{R^{\prime}}}-(\nabla_{a}u_{0})\nu^{a}\;dA=\lim_{R^{\prime}\to\infty}\int_{\partial B_{R^{\prime}}}-\frac{\partial u_{0}}{\partial r}\;d\overline{A}=\alpha(n-2)|S^{n-1}|,$
as desired.
As in Proposition 3.5 with the momentum carrier vector fields, this result now allows us to peel off a mass term and control the remainder’s decay:
Corollary 4.5.
Suppose that $g_{ab}$ satisfies Assumption 2.3 with parameters $k$, $p$, and $\tau$. Suppose that $1-n<\delta<2-n$, choose $f\in W^{k-2,p}_{\delta-2}$, and set $c=\int_{\mathbb{R}^{n}}fdV_{g}$. If $u\in W^{k,p}_{(2-n)^{+}}$ satisfies $\Delta_{g}u=f$, then
$u=cu_{0}+v,$
where $u_{0}$ is the function from Proposition 4.4 and $v\in W^{k,p}_{\delta}$.
While this reads as a statement about the Poisson equation, it can be applied to the Lichnerowicz equation (4.3) as follows: If $1-n<\tau<2-n$ and $u\in W^{k,p}_{(2-n)^{+}}$ solves (4.3) with $K_{ab}\in W^{k-1,p}_{(1-n)^{+}}$, then $\Delta_{g}u=f$ for $f\in W^{k-2,p}_{\tau-2}$ defined by
$f=\frac{1+u}{q_{n}+2}R(g)-\frac{(1+u)^{-q_{n}-1}}{q_{n}+2}|K|_{g}^{2}.$ (4.6)
Combined with Proposition 4.4(c), Corollary 4.5 now ensures that there is a constant $C$ and a function $v\in W^{k,p}_{\tau}$ such that
$\varphi:=1+u=1+\frac{C}{r^{n-2}}+v.$
This yields the expected mass term $Cr^{2-n}$ plus a remainder with prescribed decay. We observe that unlike the ADM momentum components in Proposition 3.5 and Theorem 3.9, the analogous integral $\int_{\mathbb{R}^{n}}fdV_{g}$ is not computable directly from the seed data here, so this result may not be leveraged to prescribe the ADM mass of the resulting initial data set a priori.
It remains to discuss the more delicate case of allowing threshold decay $O(r^{2-n})$ terms in $g_{ab}$, which we handle in a manner directly analogous to Section 3. Moving forward, we slightly alter Assumption 2.3 to assume that $g$ is asymptotically Euclidean of class $W^{k,p}_{(2-n)^{+}}$, still satisfying $k>1+n/p$. Given $K_{ab}\in W^{k-1,p}_{(1-n)^{+}}$, Corollary 4.2 still guarantees the existence and uniqueness of a solution $u\in W^{k,p}_{(2-n)^{+}}$ to equation (4.3) as long as $g$ is Yamabe positive, but we wish to ensure that no logarithmic terms arise and that we can control the decay of any remainder beyond the $O(r^{2-n})$ terms. We supplement Corollary 4.5 with the following (compare Proposition 3.7):
Lemma 4.6.
Suppose that $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{(2-n)^{+}}$ with $k>1+n/p$. Let $\zeta\in W^{k-2,p}(S^{n-1})$ satisfy $\int_{S^{n-1}}\zeta\;dV_{S^{n-1}}=0$, and suppose that $u\in W^{k,p}_{\delta}$ for some $\delta<0$ satisfies
$\Delta_{g}u=\zeta r^{-n}$
on the exterior region $E_{R}$ for some $R>0$. Then there is a function $v\in W^{k,p}_{(1-n)^{+}}$ and a function $\omega\in W^{k,p}(S^{n-1})$ such that, on $E_{R}$,
$u=\omega r^{2-n}+v.$
Proof.
Since $\int_{S^{n-1}}\zeta\;dV_{S^{n-1}}=0$ we can find $\omega\in W^{k,p}(S^{n-1})$ with $\Delta_{S^{n-1}}\omega=\zeta$. This function $\omega$ is unique up to an additive constant– we fix one such solution.
A straightforward computation in polar coordinates shows that $\overline{\Delta}(\omega r^{2-n})=\zeta r^{-n}$, and we define $w=\chi\omega r^{2-n}$, where $\chi$ is a cutoff function equal to $1$ outside $B_{R}$ and vanishing on $B_{R/2}$. Let p=uwp=u-w, so that on $E_{R}$, $p$ satisfies
$\Delta_{g}p=-(\Delta_{g}-\overline{\Delta})w\in W^{k-2,p}_{(2-2n)^{+}}\subset W^{k-2,p}_{(-1-n)^{+}}.$
Corollary 4.5 implies that $p=au_{0}+q$, where $a$ is a constant, $u_{0}$ is the function from Proposition 4.4, and $q\in W^{k,p}_{(1-n)^{+}}$. By the assumed decay rate for $g_{ab}$, it follows that $u_{0}$ has the form $\alpha r^{2-n}+W^{k,p}_{(1-n)^{+}}$ on $E_{R}$. We conclude that $u=p+w$ has the claimed form on $E_{R}$.
In the absence of the integral condition $\int_{S^{n-1}}\zeta\;dV_{S^{n-1}}=0$, logarithmic terms generally arise in this lemma’s conclusion. To combine this result with Corollary 4.5, we must check that the leading order $O(r^{-n})$ term in equation (4.6) coming from the scalar curvature $R(g)$ satisfies this integral condition if $g_{ab}$ is taken to have a natural form (compare to Lemma 3.8):
Lemma 4.7.
Suppose that $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{(2-n)^{+}}$ with $k>1+n/p$ and admits the decomposition
$g_{ab}=\delta_{ab}+r^{2-n}\beta_{ab}+\gamma_{ab},$ (4.7)
on $E_{R}$ for some $R>0$, where $\beta_{ab}\in W^{k,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ is homogeneous of degree zero and $\gamma_{ab}\in W^{k,p}_{\eta}$ for some $\eta<2-n$. Then the scalar curvature admits the decomposition
$R(g)=R_{\mathrm{slow}}+R_{\mathrm{fast}},$
on $E_{R}$, where $R_{\mathrm{slow}}=\zeta r^{-n}$ with $\zeta\in W^{k-2,p}(S^{n-1})$, and $R_{\mathrm{fast}}\in W^{k-2,p}_{\eta-2}\cup W^{k-2,p}_{(2-2n)^{+}}.$ Moreover, $\int_{S^{n-1}}\zeta\;dV_{S^{n-1}}=0$, and $\zeta\equiv 0$ if $\beta_{ab}$ is a multiple of $\delta_{ab}$.
Proof.
If we define $F_{ab}:=\beta_{ab}r^{2-n}$, then a computation shows that
$R_{g}=\underbrace{\frac{1}{2}\left[2\partial^{a}\partial^{b}F_{ab}-2\overline{\Delta}F^{a}_{a}\right]}_{R_{\mathrm{slow}}}+R_{\mathrm{fast}}$
where $R_{\mathrm{fast}}$ is dominated by terms of the form $\partial F\partial F$, $F\partial^{2}F$, and $\partial^{2}\gamma$, and hence lies in $W^{k-2,p}_{\rho}$ for any $\rho$ satisfying $\rho>2-2n$ and $\rho\geq\eta-2$. We note that here, and for the remainder of the proof, we perform the index operations using the metric $\delta_{ab}$.
We decompose $\beta_{ab}=m_{ab}+v_{a}n_{b}+n_{a}v_{b}+qn_{a}n_{b}$, where $n^{a}=x^{a}/r$ is the radial unit vector, where $m_{ab}$ and $v_{b}$ each annihilate $n^{a}$, and where each of $m_{ab}$, $v_{a}$, and $q$ are functions of angle only. Using the relations
$\partial_{a}r$ $=n_{a},$
$\partial_{a}n_{b}$ $=\frac{1}{r}\left[\delta_{ab}-n_{a}n_{b}\right],$
$\partial^{a}\left(\frac{qn_{a}}{r^{n-1}}\right)$ $=0,$
$\partial^{a}v_{a}$ $=\frac{1}{r}\nabla^{a}_{S^{n-1}}v_{a},$
$\partial^{a}(m_{ab})$ $=\frac{1}{r}\left[\nabla^{a}_{S^{n-1}}m_{ab}-\delta^{ac}m_{ac}n_{b}\right],$
we find that a computation shows that
$\partial^{a}\partial^{b}D_{ab}=\frac{1}{r^{n}}\left[\nabla_{S^{n-1}}^{a}\nabla_{S^{n-1}}^{b}m_{ab}+2\nabla^{a}_{S^{n-1}}v_{a}\right].$
Similarly, we have
$\overline{\Delta}D^{a}_{a}=\frac{1}{r^{n}}\Delta_{S^{n-1}}(B^{a}_{a}).$
As a consequence, we obtain
$R_{\mathrm{slow}}=\frac{1}{r^{n}}\zeta,$
where $\zeta\in W^{k-2,p}(S^{n-1})$ satisfies the condition $\int_{S^{n-1}}\zeta\;dV_{S^{n-1}}=0$.
In the case $\beta_{ab}=\delta_{ab}$ we have that $B^{a}_{a}$ is constant, $v_{a}$ is zero, and $m_{ab}$ is the metric on the sphere, and hence $\zeta=0$.
Once again, we compile the results of the present section into the following theorem summarizing the prescription of the asymptotics of the conformal factor $\varphi=1+u$ in the conformal method (compare Theorem 3.9):
Theorem 4.8.
Suppose that $g_{ab}$ satisfies Assumption 2.3 with parameters $k$, $p$, and $\tau$; suppose that $K_{ab}\in W^{k-1,p}_{\tau-1}\cup W^{k-1,p}_{(1-n)^{+}}$; and suppose that $Y_{g}>0$.
  1. (a)
    If $2-n<\tau<0$, then there exists a unique $u\in W^{k,p}_{\tau}$ for which $\varphi=1+u$ solves the Lichnerowicz equation (2.5).
  2. (b)
    If $1-n<\tau<2-n$, then the unique $u\in W^{k,p}_{(2-n)^{+}}$ for which $\varphi=1+u$ solves the Lichnerowicz equation (2.5) is given by
    $u=\frac{C}{r^{n-2}}+v$
    for some constant $C$, with the remainder $v\in W^{k,p}_{\tau}$.
  3. (c)
    In the borderline case such that $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{(2-n)^{+}}$ and has the form
    $g_{ab}=\delta_{ab}+r^{2-n}\beta_{ab}+\gamma_{ab},$ (4.8)
    on $E_{R}$ for some $R>0$, where $\beta_{ab}\in W^{k,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ is homogeneous of degree zero and $\gamma_{ab}\in W^{k,p}_{\eta}$ for some $1-n<\eta<2-n$, the unique $u\in W^{k,p}_{(2-n)^{+}}$ for which $\varphi=1+u$ solves the Lichnerowicz equation (2.5) is given by
    $u=\omega r^{2-n}+v$
    on $E_{R}$ for some $\omega\in W^{k,p}(S^{n-1})$, with the remainder $v\in W^{k,p}_{\eta}$.
Proof.
Part (a) follows immediately from Theorem 4.1.
Part (b) follows immediately from the discussion following Corollary 4.5.
We turn to part (c). The solution $u\in W^{k,p}_{(2-n)^{+}}$ to the Lichnerowicz equation (4.3), guaranteed by Corollary 4.2, satisfies $\Delta_{g}u=f$ with $f\in W^{k-2,p}_{(-n)^{+}}$ specified by equation (4.6). Here, Lemma 4.7 implies that the decomposition
$\Delta_{g}u=\zeta r^{-n}+w,$
holds on $E_{R}$, with $\zeta\in W^{k-2,p}(S^{n-1})$ satisfying $\int_{S^{n-1}}\zeta\;dV_{S^{n-1}}=0$, and with $w\in W^{k-2,p}_{\eta-2}$. Corollary 4.5 and Proposition 4.4(c) imply that the solution $u_{1}\in W^{k,p}_{(2-n)^{+}}$ to $\Delta_{g}u_{1}=w$, which exists as a consequence of Proposition 4.3, must satisfy the condition $u_{1}=Cr^{2-n}+u_{2}$ on $E_{R}$ for some constant $C$ and some $u_{2}\in W^{k,p}_{\eta}$. We now have $\Delta_{g}(u-u_{1})=\zeta r^{-n}$ on $E_{R}$, so Lemma 4.6 indicates that there is a function $\omega_{0}\in W^{k,p}(S^{n-1})$ and a function $u_{3}\in W^{k,p}_{(1-n)^{+}}$ such that $u-u_{1}=\omega_{0}r^{2-n}+u_{3}$ on $E_{R}$. Combining these results, we have on $E_{R}$ that
$u=\omega_{0}r^{2-n}+u_{1}+u_{3}=(\omega_{0}+C)r^{2-n}+u_{2}+u_{3}.$
Identifying $\omega=\omega_{0}+C$ and $v=u_{2}+u_{3}$ completes the proof.
We observe that in part (c), the resulting physical metric $\tilde{g}_{ab}=\varphi^{q_{n}-2}g_{ab}$ has the leading order $O(r^{2-n})$ term beyond $\delta_{ab}$ given by
$h_{ab}=r^{2-n}\left(\beta_{ab}+(q_{n}-2)\omega\delta_{ab}\right),$
so that a sufficient, but almost certainly not necessary, condition guaranteeing that this is not a constant multiple of $r^{2-n}\delta_{ab}$ is that $\beta_{ab}$ is not a function on $S^{2}$ multiplying $\delta_{ab}$.
5 Constructions
In Sections 3 and 4, we have established a sequence of results which guarantee that one can use the conformal method to construct general relativistic initial data sets $(\mathbb{R}^{n},\tilde{g},\widetilde{K}_{ab})$ with a number of prescribed features. These results have largely been framed as statements regarding the analytical features of the solutions to equations (2.3) and (4.3) and are not directly connected to the properties of the the ultimate initial data set. We now make formal statements collating our results into theorems, Theorems 5.1-5.3 below, summarizing precisely what can be said regarding the goal of constructing data of each of the types (CK), (B), and (A) discussed in the introduction. Each of these theorems follows readily from the results in Sections 3 and 4, so we mostly omit further argumentation. Our first summarizing theorem pertains to the construction of type (B) data. Following immediately from the well-known results stated in Theorems 3.9(a) and 4.8(a), it contains no novel content:
Theorem 5.1.
(Type (B)) We choose $2-n<\tau<0$, and we choose $k\in\mathbb{N}$ and $p\in(1,\infty)$ satisfying $k>\ell+1+n/p$ for some $\ell\in\mathbb{N}$. Suppose that $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{\tau}$ and is Yamabe positive, that $N$ is a positive function with $\delta N=N-1\in W^{k,p}_{\tau}$, and that $A_{ab}\in W^{k-1,p}_{\tau-1}$ is traceless and symmetric. The conformal constraint equations (2.5)-(2.6) then admit a unique solution for the vector field $W^{a}\in W^{k,p}_{\tau}$ and the positive function $\varphi$ with $\varphi-1\in W^{k,p}_{\tau}$. The corresponding initial data set $(\mathbb{R}^{n},\,\tilde{g},\,\widetilde{K}_{ab})$ satisfies the Einstein (vacuum, constant mean curvature) constraint equations (2.3)-(2.4) along with the type (B) asymptotic conditions:
$\tilde{g}_{ij}$ $=\delta_{ij}+o_{\ell+1}(r^{\tau}),$
$\widetilde{K}_{ij}$ $=o_{\ell}(r^{\tau-1}).$
To illustrate with $\tau=5/2-n$, one first chooses $g_{ab}$ to be Yamabe positive and asymptotically Euclidean of class $W^{k,p}_{5/2-n}$, as well as $A_{ab}\in W^{k-1,p}_{3/2-n}$, so that $Z_{a}=\operatorname{\mathrm{div}}_{g}\left(\frac{1}{2N}A_{ab}\right)\in W^{k-2,p}_{1/2-n}$. Proposition 3.1(e) ensures that one may uniquely solve equation (2.6), that is $(P_{g,N}W)_{a}=Z_{a}$, for $W^{a}\in W^{k,p}_{5/2-n}$. One now has $K_{ab}=\frac{1}{2N}(A_{ab}-(\operatorname{{\mathbb{L}}}W)_{ab})\in W^{k-1,p}_{3/2-n}$, so Theorem 4.1 ensures that we may find a conformal factor $\varphi$ solving equation (2.5) and satisfying $\varphi-1\in W^{k,p}_{5/2-n}$. For $n=3$ and for sufficiently large $k$ ($k>3+n/p$), the conformally transformed metric $\tilde{g}=\varphi^{4}g$ has the form of equation (1.3), and the physical second fundamental form $\widetilde{K}_{ij}=\varphi^{-2}K_{ij}\in W^{k,p}_{5/2-n}$ has the form (1.4), which corresponds to type (B) data.
In attempting to repeat this procedure for either type (CK) or type (A) data, equations (1.1)-(1.2) and (1.5)-(1.6), with all the falloff weights shifted down by one, one finds that $Z_{a}\in W^{k-2,p}_{-1/2-n}$, so that the desired domain of $P_{g,N}$ now has $\delta=3/2-n<2-n$, and we cannot simply invoke Proposition 3.1(e). This obstruction is not surprising: Proposition 3.1 indicates that we should expect the solution $W^{a}$ to equation (3.1) to be contained in $W^{k,p}_{(2-n)^{+}}$, ultimately corresponding to $\widetilde{K}_{ab}\in W^{k-1,p}_{(1-n)^{+}}$ and allowing an $O(r^{1-n})$ term in $\widetilde{K}_{ab}$ which gives rise to the initial data set’s linear ADM momentum. The same obstruction arises in seeking $\varphi$: The conclusion of Theorem 4.1 is limited due to the generic appearance of an $O(r^{2-n})$ term in $\varphi$, corresponding to the mass term in the metric and screening out control of any remaining decay. While the problem of extracting the $O(r^{2-n})$ term from $\varphi$ to control the remaining decay has been treated in the literature (see Proposition 1 of bieri_brill_2025 or Theorem 1.17 of bartnik1 ), previously existing results are insufficient to guarantee the emergence of type (A) data in which the mass term $h_{ij}$ in the metric is anisotropic, or to fully understand the $O(r^{1-n})$ term in the second fundamental form. As demonstrated in Theorems 5.2 and 5.3 below, we are nevertheless able to treat both of these concerns.
Our second summarizing theorem, relevant to the construction of type (CK) data, partially overcomes these hurdles, the remainder being left for our final theorem. Theorem 5.2 contains the novel result that the ADM momentum components of the constructed initial data set may always be fully anticipated, and hence prescribed as a consequence of one’s choice of the seed data, thereby allowing one to build initial data sets in the center of mass frame with nontrivial $\widetilde{K}_{ab}$ if desired. The following is a consequence of Theorem 3.9(b) (and the immediately following discussion) and Theorem 4.8(b):
Theorem 5.2.
(Type (CK)) We choose $1-n<\tau<2-n$, and we choose $k\in\mathbb{N}$ and $p\in(1,\infty)$ satisfying $k>\ell+1+n/p$ for some $\ell\in\mathbb{N}$. Suppose that $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{\tau}$ and is Yamabe positive, that $N$ is a positive function with $\delta N=N-1\in W^{k,p}_{\tau}$, and that $A_{ab}\in W^{k-1,p}_{\tau-1}$ is traceless and symmetric. The conformal constraint equations (2.5)-(2.6) then admit a unique solution for the vector field $W^{a}\in W^{k,p}_{(2-n)^{+}}$ and the positive function $\varphi$ with $\varphi-1\in W^{k,p}_{(2-n)^{+}}$. The corresponding initial data set $(\mathbb{R}^{n},\,\tilde{g},\,\widetilde{K}_{ab})$ has ADM momentum components
$p_{j}:=\mathcal{P}_{j}(\widetilde{K}_{ab})=\mathcal{G}_{j}\left(\frac{1}{2N}A_{ab}\right)=-\mathcal{R}_{j}\left(\frac{1}{2N}A_{ab}\right):=-\int_{\mathbb{R}^{n}}\nabla^{b}\left(\frac{1}{2N}A_{ab}\right)k^{a}_{(j)}\;dV_{g},$ (5.1)
and there is a constant $C$ such that
$\tilde{g}_{ij}$ $=\left(1+\frac{C}{r^{n-2}}\right)\delta_{ij}+o_{\ell+1}(r^{\tau}),$
$\widetilde{K}_{ij}$ $=\frac{2n(n-2)C_{n}}{r^{n-1}}\left[p_{i}n_{j}+p_{j}n_{i}+((n-2)n_{i}n_{j}-\delta_{ij})p_{k}n^{k}\right]+o_{\ell}(r^{\tau-1}).$
We recall that $k_{(j)}^{a}$ denotes the vector fields of Proposition 3.2, which comprise a particular basis of the kernel of $P_{g,N}$ on $W^{k,p}_{0^{+}}$. In order to use this result to prescribe the ADM momentum components in practice, one should fix $g$ and $N$ according to the above hypotheses, numerically construct the kernel vector fields $k^{a}_{(j)}$, and compute each of the integrals (5.1) for $n$ different choices of $A_{ab}$. This procedure generically produces $n$ independent momentum vectors, resulting in a basis that one can use to construct any desired momentum vector by specifying an appropriate linear combination of one’s choice of the $A_{ab}$’s. To most simply construct an initial data set in the center of mass frame, one should choose $A_{ab}=0$. To construct an initial data set in the center of mass frame which is not time-symmetric, one should compute the momentum vector for one more choice of $A_{ab}$ (for a total of $n+1$ choices, necessarily producing a linearly dependent set of momentum vectors), thereby allowing one to build a nontrivial linear combination with zero momentum.
Our third and final summarizing theorem is relevant to the construction of type (A) data. It contains the novel result that an anisotropic mass term can be built into the metric $\tilde{g}_{ij}$ of one’s initial data set in practice, providing an intermediate regime between type (B) and type (CK) data. It further indicates that a Euclidean transverse traceless tensor which is homogeneous of degree 1n1-n can be built into $\widetilde{K}_{ij}$ as its leading order term if desired– this is precisely the momentum analogue of the anisotropic mass term allowed for in $\tilde{g}_{ij}$. The following statement is a consequence of Theorem 3.9(3.25) and Theorem 4.8(c) (and the immediately following discussion):
Theorem 5.3.
(Type (A)) We choose $1-n<\tau<2-n$, and we choose $k\in\mathbb{N}$ and $p\in(1,\infty)$ satisfying $k>\ell+1+n/p$ for some $\ell\in\mathbb{N}$. Suppose that $N$ is a positive function with $\delta N=N-1\in W^{k,p}_{(2-n)^{+}}$. Suppose that $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{(2-n)^{+}}$, is Yamabe positive, and is given by
$g_{ij}=\delta_{ij}+r^{2-n}\beta_{ij}+\gamma_{ij}$
on $E_{R}$ for some $R>0$, where $\beta_{ij}\in W^{k,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ is homogeneous of degree zero and $\gamma_{ij}\in W^{k,p}_{\tau}$. Suppose that $A_{ab}\in W^{k-1,p}_{(1-n)^{+}}$ has the form
$A_{ab}=r^{1-n}B_{ab}+C_{ab}$ (5.2)
on $E_{R}$, where $B_{ab}\in W^{k-1,p}_{\mathrm{loc}}(\mathbb{R}^{n}\setminus\{0\})$ is homogeneous of degree zero and $C_{ab}\in W^{k-1,p}_{\tau-1}$, both traceless and symmetric. The conformal constraint equations (2.5)-(2.6) admit a unique solution for the vector field $W^{a}\in W^{k,p}_{(2-n)^{+}}$ and the positive function $\varphi$ with $\varphi-1\in W^{k,p}_{(2-n)^{+}}$. The corresponding initial data set $(\mathbb{R}^{n},\,\tilde{g},\,\widetilde{K}_{ab})$ has ADM momentum components
$\mathcal{P}_{j}(\widetilde{K}_{ab})=\mathcal{G}_{j}\left(\frac{1}{2N}A_{ab}\right):=\int_{\mathbb{R}^{n}}\frac{1}{2N}A_{ab}\nabla^{b}k^{a}_{(j)}\;dV_{g}.$
There is a homogeneous of degree 2n2-n symmetric tensor $h_{ij}$ and a homogeneous of degree 1n1-n Euclidean transverse traceless tensor $D_{ij}$ such that
$\tilde{g}_{ij}$ $=\delta_{ij}+h_{ij}+o_{\ell+1}(r^{\tau}),$
$\widetilde{K}_{ij}$ $=D_{ij}+o_{\ell}(r^{\tau-1}).$
Moreover: if $B_{ab}\neq\omega\delta_{ab}$ for any $\omega\in W^{k,p}(S^{n-1})$, then $h_{ij}$ is not a constant multiple of $r^{2-n}\delta_{ij}$.
We remark that the discussion at the end of Section 3 further implies that, with an appropriate choice of $A_{ab}\in W^{k-1,p}_{(1-n)^{+}}$, one may arrange that the leading order term in $\widetilde{K}_{ij}$ can be any given Euclidean transverse traceless tensor $D_{ij}$ that is homogeneous of degree 1n1-n. That is, this term can be prescribed, and it could readily be made nontrivial even with zero ADM momentum.
Taken together, Theorems 5.1-5.3 answer the question of how one can practically construct initial data sets of each of the types (CK), (B), and (A) (and a bit beyond) using the formalism of the conformal method. These theorems indicate that one can precisely control the initial data set’s ADM momentum and the leading (or next-to-leading) order decay rates of the physical metric $\tilde{g}_{ij}$ and the second fundamental form $\widetilde{K}_{ij}$, and that one may build an anisotropic mass term into $\tilde{g}_{ij}$, or the analogous term into $\widetilde{K}_{ij}$, if desired. As an application, these results guarantee that a simple example numerically constructed below is of type (A), allowing us to place the strongest known conditions (cf. bieri_brill_2025 ) on the collection of spacetimes to which the antipodal matching conjecture of strominger could apply.
6 A Numerical Example
In this section, we present a numerically computed Brill wave solution of type (A) using the method of bieri_brill_2025 . Brill waves are initial data sets on $\mathbb{R}^{3}$ for which $\widetilde{K}=0$ and $\tilde{g}$ is axisymmetric. The momentum constraint is then trivially satisfied, while the Lichnerowicz equation (2.5) becomes
$8{\Delta_{g}}\varphi+R(g)\varphi=0.$ (6.1)
We find it convenient to perform the numerical computations in spherical polar coordinates $(r,\theta,\phi)$. In these coordinates, the line element of the unphysical metric $g$ takes the form
$d{s^{2}}={e^{2q}}(d{r^{2}}+{r^{2}}d{\theta^{2}})+{r^{2}}{\sin^{2}}\theta d{\phi^{2}}.$ (6.2)
Here, in keeping with axisymmetry, the function $q$ depends only on $r$ and $\theta$. Equation (6.1) then takes the form
$\overline{\Delta}\varphi+S\varphi=0,$ (6.3)
where $\overline{\Delta}$ is the flat Laplacian and $S$ is given by
$S=-{\frac{1}{4}}\left({\frac{{\partial^{2}}q}{\partial{r^{2}}}}+{\frac{1}{r}}\,{\frac{\partial q}{\partial r}}+{\frac{1}{r^{2}}}\,{\frac{{\partial^{2}}q}{\partial{\theta^{2}}}}\right).$ (6.4)
We define $F:=\ln\varphi$ and decompose $F={F_{1}}+{F_{2}}$, with $F_{1}$ a solution to
$\overline{\Delta}{F_{1}}=S.$ (6.5)
Equation (6.3) now becomes
$\overline{\Delta}{F_{2}}=-{\vec{\nabla}}F\cdot{\vec{\nabla}}F,$ (6.6)
which we seek to solve for $F_{2}$. Given a numerical method to invert the flat Laplacian, we first solve equation (6.5) and proceed to solve equation (6.6) by iteration. That is, having found $F_{1}$, we make the initial guess of zero for $F_{2}$ and repeatedly solve equation (6.6) for an improved version of $F_{2}$, where the right hand side of equation (6.6) is computed using the previous version of $F_{2}$. At each step, the current version of $F_{2}$ is stored, with the final version used to compute $F$.
Our numerical method to invert the flat Laplacian relies on the standard Green’s function for axisymmetric functions (see e.g., Jackson ) where we compute all integrals numerically. Explicitly, for equation (6.5) we have
${F_{1}}={\sum_{\ell=0}^{\infty}}\left[{g_{\ell}}(r){r^{-(\ell+1)}}+{h_{\ell}}(r){r^{\ell}}\right]{P_{\ell}}(\cos\theta),$ (6.7)
where ${g_{\ell}}(r)$ and ${h_{\ell}}(r)$ are given by
${g_{\ell}}(r)={\int_{0}^{r}}{{\tilde{r}}^{\ell+2}}d{\tilde{r}}{\int_{0}^{\pi}}\sin\theta\,d\theta\,S({\tilde{r}},\theta){P_{\ell}}(\cos\theta),$ (6.8)
${h_{\ell}}(r)={\int_{r}^{\infty}}{{\tilde{r}}^{1-\ell}}d{\tilde{r}}{\int_{0}^{\pi}}\sin\theta\,d\theta\,S({\tilde{r}},\theta){P_{\ell}}(\cos\theta),$ (6.9)
and $P_{\ell}$ are the Legendre polynomials.
To assess whether our constructed initial data set conforms to the antipodal conjecture of strominger , we examine the behavior of the curvature component $\rho$, which is the principal curvature corresponding to the plane that contains the gradient of $r$ and the normal to the initial data surface. For Brill wave initial data, $\rho$ is given asymptotically (as $r\to\infty$) by the following expression:
${r^{2}}\rho$ $=$ $-{r^{2}}{\frac{{\partial^{2}}q}{\partial{r^{2}}}}-{\frac{{\partial^{2}}q}{\partial{\theta^{2}}}}-\cot\theta{\frac{\partial q}{\partial\theta}}+2{r^{2}}{\frac{\partial q}{\partial r}}{\frac{\partial F}{\partial r}}-2{\frac{\partial q}{\partial\theta}}{\frac{\partial F}{\partial\theta}}$ (6.10)
$-4{r^{2}}{\frac{{\partial^{2}}F}{\partial{r^{2}}}}-4r{\frac{\partial F}{\partial r}}-2{\frac{{\partial^{2}}F}{\partial{\theta^{2}}}}-2\cot\theta{\frac{\partial F}{\partial\theta}}-4{{\left({\frac{\partial F}{\partial\theta}}\right)}^{2}}.$
Figure 1: $\varphi$ with $q$ given by equation (6.11) with ${a_{0}}=1,\,{r_{0}}=10,\,\gamma=2$
In order to produce an example that is not antipodally symmetric, we consider $q$ of the form
$q={a_{0}}{r^{3}}\cos(\theta){\sin^{2}}(\theta){{({r^{2}}+{r_{0}^{2}})}^{-\gamma}},$ (6.11)
where ${a_{0}},\,{r_{0}},\,$ and $\gamma$ are constants. In order to obtain type (A) behavior, we choose $\gamma=2$– that the resulting initial data set is of type (A) is now implied by Theorem 5.3. Figure 1 graphs the numerically computed $\varphi$ for the case ${a_{0}}=1,{r_{0}}=10,\gamma=2$ up to the radius of $r=100$. Figure 2 plots ${r^{3}}\rho$ for two different values of $r$: $r=10000$ and $r=15000$. The fact that the two curves agree corroborates the claim that the asymptotic behavior of $\rho$ is $\rho\propto{r^{-3}}$, as expected.
Figure 2: ${r^{3}}\rho$ as a function of $\theta$ for $r=10000$ and $r=15000$ with $q$ given by equation (6.11) with ${a_{0}}=1,\,{r_{0}}=10,\,\gamma=2$
We observe in Figure 2 that $\rho$ is not antipodally symmetric. This is relevant to the conjecture of Strominger strominger that in the limit of early time at null infinity there is a symmetry involving the combination of time reflection and antipodal mapping. Since Brill wave initial data sets have zero extrinsic curvature, it follows that their time evolution has time reflection symmetry. The evolution of the Brill wave initial data set whose features are plotted in Figures 1 and 2 therefore cannot satisfy the conjecture of strominger . This result should not be regarded as a counterexample to this conjecture, but rather as a statement regarding the sort of asymptotic flatness under which this conjecture can hold: that is, the conjecture of strominger does not hold for metrics of type (A).
7 Conclusion
In this work, we have thoroughly investigated the problem of constructing maximal solutions $(\mathbb{R}^{n},\tilde{g},\widetilde{K}_{ab})$ to the vacuum constraint equations (2.3)-(2.4) via the conformal method with prescribed asymptotic behavior. While it has long been known that one can construct unique solutions for any choice of seed data $g_{ab}$, $A_{ab}$, and $N$ taken in appropriate function spaces, we have clarified the precise extent of the technical control which can be exerted over the asymptotics of these solutions. At a broad level, we have proceeded by first characterizing the solutions to the conformal constraint equations (2.5)-(2.6) explicitly at leading order, using the “momentum carrier” vector fields $W^{a}_{(j)}$ and the “mass carrier” function $u_{0}$ in the event of rapid decay in the seed data, or using understood solutions from analogous problems involving the flat operators $\overline{P}$ and $\overline{\Delta}$ in the event of threshold decay corresponding to the natural forms (3.24) and (4.8) for $A_{ab}$ and $g_{ab}$. Generally, we have shown that peeling off this leading-order behavior of the solutions renders the equations solvable in a space with faster decay, asserting control of the decay rate of subleading corrections. Theorems 3.9 and 4.8 summarize these results at the more abstract level of the vector field $W^{a}$ and conformal factor $\varphi$ directly solved for in the conformal method, while Theorems 5.1, 5.2, and 5.3 summarize these results at the level of the ultimately constructed initial data set $(\mathbb{R}^{n},\tilde{g},\widetilde{K}_{ab})$. We remark that the leading-order terms may be understood well enough to prescribe the initial data set’s ADM momentum components, and (in working with prescribed threshold decay) to impose that the metric include an anisotropic mass term.
The results established herein ensure that, with an appropriate choice of seed data, one can construct initial data sets of any of the three types discussed in the introduction. We have illustrated this in Section 6 by numerically constructing a simple example of a type (A) spacetime guaranteed to include an anisotropic mass term, which previously available results were unable to handle. This example is constructed such that its spacetime evolution cannot satisfy the antipodal matching conjecture of strominger , restricting the collection of spacetimes within which one can expect this property of physical interest to be realized. It is our hope that our results furnish the ability to further construct a wealth of initial data sets with various decay features, allowing the relativity community to probe, both numerically and analytically, how physically interesting properties of spacetimes may depend on one’s choice of physically motivated asymptotic behavior.
Acknowledgements
Lydia Bieri is supported by NSF grant DMS-2204182 to The University of Michigan. James Wheeler was also partially supported by NSF grant DMS-2204182 to The University of Michigan. David Garfinkle was supported by NSF grant PHY-2102914 to Oakland University.
Appendix A Fredholm Properties of the Vector Laplacian
This appendix contains auxiliary technical results on the Fredholm properties of vector Laplacians. Although the body of our work above is restricted to the setting of metrics with Hölder continuous first derivatives, for the sake of generality and potential other applications we now relax the condition $k>n/p+1$ of Assumption 2.3 with $k>n/p$ instead.
Assumption A.1.
The metric $g_{ab}$ on $\mathbb{R}^{n}$, $n\geq 3$ and the lapse $N$ satisfy the following two conditions for some $\tau<0$, $k\in\mathbb{N}$, and $1<p<\infty$ with $k>n/p$:
  • $g_{ab}$ is asymptotically Euclidean of class $W^{k,p}_{\tau}$.
  • NN is a positive function with $\delta N=N-1\in W^{k,p}_{\tau}$
The vector Laplacian is then $P_{g,N}=\operatorname{\mathrm{div}}_{g}(\frac{1}{2N}\operatorname{{\mathbb{L}}})$, which we abbreviate as $P$ in this section. The model Euclidean vector Laplacian with $N=1$ is then $\overline{P}=\frac{1}{2}\overline{\operatorname{\mathrm{div}}}(\overline{\operatorname{{\mathbb{L}}}})$.
The natural Fredholm theory for a self-adjoint operator such as the vector Laplacian makes use of dual function spaces, and in particular distributions having a negative order of differentiability. Hence we generalize Definition 2.1 to allow weighted Sobolev spaces $W^{k,p}_{\delta}$ with $k\leq 0$. Recall that if $k<0$, then $W^{k,p}(\Omega)$ for an arbitrary domain $\Omega\subset\mathbb{R}^{n}$ is the topological dual space $W^{k,p}(\Omega):=(W^{-k,p^{\prime}}_{0}(\Omega))^{*}$, with the induced dual norm, where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$ and where $W^{-k,p^{\prime}}_{0}(\Omega)$ is the closure of $C_{c}^{\infty}(\Omega)$ in $W^{-k,p^{\prime}}(\Omega)$.
Definition A.2.
Let $k\in\mathbb{Z}$, $1<p<\infty$, and $\delta\in\mathbb{R}$. We denote by $A_{r}$ the annulus $\{x\in\mathbb{R}^{n}:r/2<|x|<2r\}$, and we let $S_{r}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be the scaling map $S_{r}(x)=rx$, so $S_{r}(A_{1})=A_{r}$. The weighted Sobolev space $W^{k,p}_{\delta}$ consists of the tempered distributions $u$ on $\mathbb{R}^{n}$ such that
[…]
It follows from Bartnik’s scaling technique bartnik1 that for $k\geq 0$ the norm above is equivalent to that of Definition 2.1. Although we do not require it, we remark that Definition A.2 generalizes to non-integral scales of differentiabilty, and indeed elementary arguments show that $W^{k,p}_{\delta}$ defined here agrees with the space $f^{s}_{p,2,\mu}$ of triebel_spaces_1976a with $s=k$ and $\mu=-\delta p+sp-n$. See also the analogous definition for the special case $p=2$ in maxwell_rough_2006 . General properties of these spaces in the case $k\geq 0$ can be found in bartnik1 , and their generalizations to the broader family appear in triebel_spaces_1976a and maxwell_rough_2006 . The sole exception is the compactness of the embedding $W^{k,p}_{\delta}\hookrightarrow W^{j,p}_{\delta^{\prime}}$ when $k\geq j$ and $\delta^{\prime}<\delta$. This is appears in the special case $p=2$ in maxwell_rough_2006 and can be extended to general $p$ using the same technique.
The following multiplication lemma provides criteria under which the product of elements of two weighted Sobolev spaces determines an element of a third weighted space. The proof follows from Bartnik’s scaling technique and the corresponding Sobolev multiplication properties on unweighted spaces (see, for example, Theorem 2.5 of holst_scaling_2023 ). See also allen_sobolevclass_2022a Proposition 3.11, which is the equivalent result for asymptotically hyperbolic manifolds.
Lemma A.3.
Suppose $1<p_{1},p_{2},q<\infty$, $k_{1},k_{2},j\in\mathbb{Z}$ and $\delta_{1},\delta_{2}\in\mathbb{R}$. Let $r_{1},r_{2}$ and $r$ be defined by
$\frac{1}{r_{i}}=\frac{1}{p_{i}}-\frac{k_{i}}{n}\quad i=1,2\quad\text{and}\quad\frac{1}{r}=\frac{1}{q}-\frac{j}{n}.$
Pointwise multiplication of $C^{\infty}_{c}$ functions extends to a continuous bilinear map $W^{k_{1},p_{1}}_{\delta_{1}}\times W^{k_{2},p_{2}}_{\delta_{2}}\rightarrow W^{j,q}_{\delta_{1}+\delta_{2}}(\mathbb{R})$ so long as
$\frac{1}{p_{1}}+\frac{1}{p_{2}}$ $\geq\frac{1}{q},$
$k_{1}+k_{2}$ $\geq 0,$
$\min({k_{1},k_{2}})$ $\geq j,$
$\max\left(\frac{1}{r_{1}},\frac{1}{r_{2}}\right)$ $\leq\frac{1}{r},$
$\text{and}\qquad\quad\frac{1}{r_{1}}+\frac{1}{r_{2}}$ $\leq\min\left(1,\frac{1}{r}\right),$
with the final inequality being strict if $\min(1/r_{1},1/r_{2},1-1/r)=0$.
A straightforward application of Lemma A.3 (see Proposition 2.6 of holst_scaling_2023 for the corresponding result in unweighed spaces) establishes the following elementary mapping property of the vector Laplacian:
$P:W^{j,q}_{\delta}\to W^{j-2,q}_{\delta-2}$
is continuous so long as
$2-k$ $\leq j\leq k,$ (A.1)
$\frac{1}{p}-\frac{k}{n}$ $\leq\frac{1}{q}-\frac{j}{n}\leq\frac{1}{p^{\prime}}-\frac{2-k}{n}.$ (A.2)
Definition A.4.
Suppose $1<p<\infty$ and $k>n/p$. The set $\mathcal{S}^{k,p}$ of compatible Sobolev indices is the collection of pairs $(j,q)$ in $\mathbb{Z}\times(1,\infty)$ satisfying A.1A.2.
Using the standing assumption $k>n/p$ it is easy to show $\mathcal{S}^{k,p}$ is nonempty and contains $j=1$, $p=2$. This observation and related facts concerning $\mathcal{S}^{k,p}$ are discussed in holst_scaling_2023 .
The dual space $(W^{j,q}_{\delta})^{*}$ for scalars can be identified with $W^{-j,q^{\prime}}_{-n-\delta}$; see triebel_spaces_1976a for one approach. This identification does not rely on a metric, but it is more useful for our purposes to use a metric-specific representation of the dual space that aligns with the adjoint structure of $P$. The following result is a variation of Proposition 3.18 of allen_sobolevclass_2022a , which proves the analogous fact in the asymptotically hyperbolic setting. The proof follows the same technique, and is omitted.
Lemma A.5.
Suppose $g_{ab}$ satisfies Assumption A.1 with the parameters $k$, $p$, and $\tau$. Suppose that $\delta\in\mathbb{R}$, and that $(j,q)\in\mathbb{Z}\times(1,\infty)$ satisfies $|j|\leq k$ and
$\frac{1}{p}-\frac{k}{n}\leq\frac{1}{q}-\frac{j}{n}\leq\frac{1}{p^{\prime}}+\frac{k}{n}.$ (A.3)
Given $X\in W^{-j,q^{\prime}}_{-n-\delta}$, we define $f_{X}:C^{\infty}_{c}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ by
$f_{X}(Y):=\int_{\mathbb{R}^{n}}\left<X,Y\right>_{g}\;dV_{g}.$ (A.4)
Then $f_{X}$ extends to a continuous map
$f_{X}:W^{j,q}_{\delta}\to\mathbb{R}.$
Moreover the map $X\mapsto f_{X}$ is a linear isomorphism $W^{-j,q^{\prime}}_{-n-\delta}\rightarrow(W^{j,q}_{\delta})^{*}$.
Using the specific identification of $(W^{j,q}_{\delta})^{*}$ with $W^{-j,q^{\prime}}_{-n-\delta}$ from Lemma A.5, the following lemma shows that the vector Laplacian is self-adjoint.
Lemma A.6.
Suppose $g_{ab}$ satisfies Assumption A.1 with the parameters $k$, $p$, and $\tau$. Consider Sobolov parameters $(j,q)$ and a weight $\delta$ such that $(j,q)\in\mathcal{S}^{k,p}$ from Definition A.4. Then $(-j,q^{\prime})\in S^{k,p}$ as well, and hence $P$ acts continuously on $W^{-j,p^{\prime}}_{2-n-\delta}$. Moreover, for all $X\in W^{j,q}_{\delta}$ and $Y\in W^{-j,q^{\prime}}_{2-n-\delta}$,
$\int_{\mathbb{R}^{n}}\left<PX,Y\right>_{g}\;dV_{g}=\int_{\mathbb{R}^{n}}\left<X,PY\right>_{g}\;dV_{g}.$ (A.5)
Proof.
The fact that $(-j,q^{\prime})\in\mathcal{S}^{k,p}$ is an easy computation from the definition. Equality (A.5) holds when $X$ and $Y$ are smooth and compactly supported. The continuity of $P$ on $W^{j,q}_{\delta}$ and $W^{-j,q^{\prime}}_{2-n-\delta}$ together with Lemma A.5 and the density of smooth compactly supported vector fields in these spaces establishes equality (A.5) generally.
We have the following elementary elliptic regularity estimate, a variation of bartnik1 Proposition 1.6.
Lemma A.7.
Suppose that $g_{ab}$ and $N$ satisfy Assumption A.1 with parameters $k$, $p$, and $\tau$. Assume $X\in W^{2-k,q^{\prime}}_{\mathrm{loc}}$ so that $PX$ is well-defined as a distribution. Suppose for some $(j,q)\in\mathcal{S}^{k,p}$ from Definition A.4 and some $\delta\in\mathbb{R}$ that $X\in W^{j-2,q}_{\delta}$ and that $PX=Z$ for some $Z\in W^{j-2,q}_{\delta-2}$. Then $X\in W^{j,q}_{\delta}$ and
[…]
Proof.
Theorem 2.21 of holst_scaling_2023 implies $X\in W^{j,q}_{\mathrm{loc}}$. To see that $X$ additionally lies in the weighted space we first decompose
$P=\overline{P}+\sum_{|\alpha|\leq 2}A^{\alpha}\partial_{\alpha}$
with matrix coefficients $A^{\alpha}\in W^{k-2+|\alpha|,p}_{\tau-2+|\alpha|}$. Recalling the scaling operator $S_{r}$ of Definition A.2 we find that for any $r=2^{m}$ for some $m\in\mathbb{N}$
$S_{r}^{*}(PX)=r^{-2}\overline{P}S_{r}^{*}(X)+\sum_{|\alpha|\leq 2}r^{-|\alpha|}S_{r}^{*}(A^{\alpha})\partial_{\alpha}S_{r}^{*}X.$
Hence
$\overline{P}S_{r}^{*}(X)=r^{2}S_{r}^{*}(PX)-r^{\tau}\sum_{|\alpha|\leq 2}r^{2-|\alpha|-\tau}S_{r}^{*}(A^{\alpha})\partial_{\alpha}S_{r}^{*}X.$
Theorem 2.5 of holst_scaling_2023 on Sobolev multiplication and the embedding $\ell^{q}\hookrightarrow\ell^{\infty}$ imply
[…]
Hence elliptic estimates for $\overline{P}$ yield
[…]
where $\hat{A}^{1}$ is the thickened annulus $1/3<|x|<3$. Since the the implicit constant is independent of $r=2^{m}$ we can find $M$ sufficiently large so that if $m\geq M$ then the term multiplied by $r^{\tau}$ can be absorbed into the left hand side to find
[…]
Now
[…]
Hence multiplying inequality (A.7) by $2^{-\delta m}$, raising to the $q^{\rm th}$ power, and summing we find
[…]
Therefore
[…]
and we conclude $X\in W^{j,q}_{\delta}$. Estimate (A.6) follows from estimate (A.8) along with local regularity estimates for the finitely many terms omitted from the sum.
Fredholm properties of the vector Laplacian follow from those of the Euclidean model operator $\overline{P}$, which we establish now.
Proposition A.8.
Suppose $\delta$ is non-exceptional (i.e. $\delta$ is not an integer or $2-n<\delta<0$). Then
$\overline{P}^{j,q}_{\delta}:W^{j,q}_{\delta}\rightarrow W^{j-2,q}_{\delta-2}$
is continuous and Fredholm for all $j\in\mathbb{Z}$ and each $1<q<\infty$. Moreover:
  1. (a)
    $\ker\overline{P^{j,q}_{\delta}}$ is independent of $j$ and $q$ and consists of polynomials. Consequently it is trivial if $\delta<0$,
  2. (b)
    $\operatorname{\mathrm{im}}\overline{P}^{j,q}_{\delta}=\left(\ker\overline{P}^{2-j,q^{\prime}}_{2-n-\delta}\right)^{\perp}$ in the following sense: a vector $Z\in\operatorname{\mathrm{im}}\overline{P}^{j,q}_{\delta}$ if and only if $\int_{\mathbb{R}^{n}}\left<Z,K\right>_{g}\;dV_{g}=0$ for all $K\in\ker\overline{P}^{2-j,q^{\prime}}_{2-n-\delta}$,
  3. (c)
    the Fredholm index $\iota\left(\overline{P}^{j,q}_{\delta}\right)$ is independent of $j$ and $q$,
  4. (d)
    if $2-n<\delta<0$, then $\overline{P}^{j,q}_{\delta}$ is an isomorphism,
  5. (e)
    for all $X\in W^{j,q}_{\delta}$ and each $\delta^{\prime}\in\mathbb{R}$
[…]
Proof.
Theorem 3 of lockhart_elliptic_1983 implies that if $j\geq 2$ and if $\delta$ is non-exceptional, then $\overline{P}^{j,q}_{\delta}$ is Fredholm. Lemma A.6 shows that $\left(\overline{P}^{j,q}_{\delta}\right)^{*}=\overline{P}^{2-j,q^{\prime}}_{2-n-\delta}$. Noting that $2-n-\delta$ is non-exceptional if and only if $\delta$ is, we conclude that $\overline{P}^{j,q}_{\delta}$ is Fredholm for $j\leq 0$ as well since the adjoint of a Fredholm map is Fredholm.
Theorem 3 of lockhart_elliptic_1983 also implies that $\ker\overline{P}^{j,q}_{\delta}$ consists of polynomials if $j\geq 2$. Lemma A.7 implies this same fact is true for fixed $q$ but arbitrary $j$. Moreover, A polynomial is in $W^{j,q}_{\delta}$ if and only if its order is less than $\delta$ and hence the kernel depends on $\delta$ but is independent of $j$ and $q$. We denote this common kernel by $\ker\overline{P}_{\delta}$.
To handle the marginal case $j=1$ we first observe from Lemma A.7 that $\ker\overline{P}^{j,q}_{\delta}$ is independent of $j$ and $q$. Since $\ker\overline{P}^{1,q}_{\delta}=\ker\overline{P}_{\delta}$ it is finite dimensional and elementary arguments imply $\operatorname{\mathrm{im}}\overline{P}^{1,q}_{\delta}\subset W^{-1,q}_{\delta-2}\cap(\ker\overline{P}_{2-n-\delta})^{\perp}$. If we show the reverse containment we have established that $\overline{P}^{1,q}_{\delta}$ has closed range and finite dimensional cokernel and hence $\overline{P}^{1,q}_{\delta}$ is Fredholm. To establish the reverse containment, suppose $Z\in W^{-1,q}_{\delta-2}\cap(\ker\overline{P}^{-1,q^{\prime}}_{-n-\delta})^{\perp}$. From our established results for $\overline{P}^{-2,q}_{\delta}$ there exists $X\in W^{0,q}_{\delta}$ with $\overline{P}X=Z$ and Lemma A.7 implies $X\in W^{1,q}_{\delta}$ as needed.
Since the image of $\overline{P}^{j,q}_{\delta}$ is characterized in terms of the kernel of the adjoint for a Fredholm map (this is item (b)), we see that the dimension of the cokernel is independent of $j$ and $q$, and hence so is the index.
The isomorphism range $2-n<\delta<0$ follows from the fact that for these values of $\delta$ we have $\delta<0$ and $2-n-\delta<0$, so the kernels of both $\overline{P}^{j,q}_{\delta}$ and its adjoint $\overline{P}^{2-j,q^{\prime}}_{2-n-\delta}$ are trivial.
Finally, inequality (A.9) is a special case of a more general fact about Fredholm operators. If $T:X\rightarrow Y$ is a semi-Fredholm map between Banach spaces, and if $Q:X\to Z$ is a continuous map that is injective on $\ker T$, then for all $x\in X$,
[…]
The next two results concern the asymptotics of solutions of $\overline{P}X=Z$. The first shows that if $Z$ has $O(r^{\delta-2})$ decay then $\overline{P}X=Z$ is solvable, up to an error on a compact set, with $X$ having $O(r^{\delta})$ decay. The second concerns the specific asymptotic structure of a solution of $PX=Z$ when $Z$ has compact support.
Lemma A.9.
Suppose $k\in\mathbb{Z}$, $1<p<\infty$ and that $\delta$ is non-exceptional. If $Z_{a}\in W^{k-2,p}_{\delta-2}$, then for any fixed $R>0$ there exists $X^{a}\in W^{k,p}_{\delta}$ such that $(\overline{P}X)_{a}=Z_{a}$ on $E_{R}$.
Proof.
We need only consider the case $\delta<2-n$, since $\overline{P}$ is surjective otherwise.
Recall that the adjoint of $\overline{P}^{k,p}_{\delta}$ is $\overline{P}^{2-k,p^{\prime}}_{2-n-\delta}$. Letting $\{H^{a}_{(j)}\}_{j=1}^{m}$ be a basis for $\ker\overline{P}^{2-k,p^{\prime}}_{2-n-\delta}$, Proposition A.8 implies that the components of each $H^{a}_{(j)}$ are polynomials.
Define $T:C^{\infty}_{c}(B_{R})\to\mathbb{R}^{m}$ via $T(X)_{j}=\int_{B_{R}}H_{(j)}^{a}X_{a}\;dV$. This map is surjective, for otherwise there would exist a nonzero $\beta\in\mathbb{R}^{m}$ with $\beta\cdot T(X)=0$ for all compactly supported smooth vector fields on $B_{R}$. As a consequence $\beta^{j}H^{a}_{(j)}$ vanishes on $B_{R}$. But since the elements of the kernel are polynomials, we find that $\beta^{j}H^{a}_{(j)}=0$ on $\mathbb{R}^{n}$, contradicting the linear independence of the basis elements.
Now fix $Z_{a}\in W^{k-2,p}_{\delta-2}$. From the surjectivity of $T$ we can find $Y_{a}\in C^{\infty}_{c}(B_{R})$ with $T(Y)_{j}=\int_{\mathbb{R}^{n}}H^{a}_{(j)}Z_{a}\;d\overline{V}$ for each $j$. Then $Z_{a}-Y_{a}$ is $L^{2}$ orthogonal to the kernel of the adjoint of $\overline{P}^{k,p}_{\delta}$ and Proposition A.8 implies that there exists $X^{a}\in W^{k,p}_{\delta}$ solving $(\overline{P}X)_{a}=Z_{a}-Y_{a}$. Since $Y_{a}=0$ outside $B_{R}$, the proof is complete.
We now establish multipole expansions for solutions of $\overline{P}X=Z$ when $Z$ has compact support. The monopole term in this expansion is used to construct the momentum carrier vector fields $W^{a}_{(j)}$ of Proposition 3.4.
Lemma A.10.
Suppose $k\in\mathbb{Z}$, $1<p<\infty$, and that $\delta<0$ is non-exceptional. Let $X^{a}\in W^{k,p}_{\delta}$ be a vector field such that $(\overline{P}X)_{a}$ is supported on $B_{R}$ for some $R>0$ and fix $\ell\in\mathbb{N}$. There exist constants $c_{b}^{\alpha}$ (depending on indices $1\leq b\leq n$ and multiindices $\alpha$ with $|\alpha|\leq\ell$) along with a vector field $Y^{a}\in W^{k,p}_{(2-n-\ell)^{+}}$ such that on $E_{2R}$
$X^{a}=\sum_{|\alpha|\leq\ell}c_{b}^{\alpha}\partial_{\alpha}G^{ab}+Y^{a}.$ (A.10)
Here, $G^{ab}$ is the Green’s function from equation (3.10).
Proof.
Let $Z_{a}=(\overline{P}X)_{a}$ and let
$U^{a}(x)=\int G^{ab}(x-y)Z_{b}(y)\;d\overline{V}(y)$
with the convolution meant in the sense of distributions if $k<2$. Then $\overline{P}U=Z$ and the proof proceeds by showing first that $U^{a}$ admits the decomposition (A.10), and then that $U^{a}=X^{a}$ to complete the proof.
For each multiindex $\alpha$ with $|\alpha|\leq\ell$, define $c_{b}^{\alpha}:=\int\frac{(-y)^{\alpha}}{\alpha!}Z_{b}(y)\;d\overline{V}(y)$; this quantity is well defined since $Z_{b}$ has compact support. Note that if $\alpha=(\alpha_{1},\ldots,\alpha_{n})$ then $z^{\alpha}=z_{1}^{\alpha_{1}}\cdots z_{n}^{\alpha_{n}}$ and $\alpha!=\alpha_{1}!\cdots\alpha_{n}!$. We then write
$\int G^{ab}(x-y)Z_{b}(y)\;d\overline{V}(y)=\sum_{|\alpha|\leq\ell}c_{b}^{\alpha}\partial_{\alpha}G^{ab}(x)+W^{a}$
with
$W^{a}(x)=\int_{\mathbb{R}^{n}}H^{ab}(x,y)F_{b}(y)\;d\overline{V}(y),$
and with
$H^{ab}(x,y)=G^{ab}(x-y)-\sum_{|\alpha|\leq\ell}\frac{(-y)^{\alpha}}{\alpha!}\partial_{\alpha}G(x-y).$
On the region $|x|\geq 2R$, we find that $W^{a}$ is smooth since it is the convolution of a smooth function with $Z_{a}$. Since $G^{ab}$ is smooth and homogeneous of degree 2n{2-n} on $\mathbb{R}^{n}\setminus\{0\}$, if $|y|\leq R$ and if $|x|\geq 2R$ then
$|H^{ab}(x,y)|\lesssim|x|^{2-n-\ell}|y|^{\ell}.$
For each multi-index $\alpha$, the derivative $\partial^{\alpha}G^{ab}$ is homogeneous of order $2-n-|\alpha|$, and the same argument then shows that
$|\partial^{\alpha}_{x}(H^{ab}(x,y))|\lesssim|x|^{2-n-|\alpha|-\ell}|y|^{\ell}.$
Hence on the region $|x|\geq 2R$ we have uniform estimates $|\partial^{\alpha}W^{a}|\lesssim|x|^{2-n-|\alpha|-\ell}$.
Let $\chi$ be a cutoff function equal to $1$ on $B_{R}$ and vanishing on $E_{2R}$. We then have
$U^{a}=(1-\chi)\left[\sum_{|\alpha|\leq\ell}c^{\alpha}_{b}\partial_{\alpha}G^{ab}\right]+\underbrace{\chi U^{a}+(1-\chi)W^{a}}_{=:Y^{a}}.$
The uniform decay estimates on the derivatives of $W^{b}$ and the compact support of $\chi Z^{a}$ show that $Y^{a}$ belongs to $W^{k,p}_{(2-n-\ell)^{+}}$. Hence $U^{a}$ admits the decomposition (A.10), and that same decomposition implies $U^{a}\in W^{k,p}_{(2-n)^{+}}$. Since $\delta<0$ and since $(\overline{P}X)_{a}=(\overline{P}Z)_{a}$, we conclude $X^{a}=Z^{a}$ to complete the proof.
We now return to the general case of the vector Laplacian $P=P_{g,N}$. A standard perturbation technique (compare with bartnik1 Theorem 1.10) provides the following coercivity estimate, generalizing equation (A.9):
Proposition A.11.
Suppose that $g_{ab}$ and $N$ satisfy Assumption A.1 with parameters $k$, $p$, and $\tau$. Suppose $(j,q)\in\mathcal{S}^{k,p}$, and that $\delta$ is non-exceptional. Then for all $X\in W^{j,q}_{\delta}$ and any $\delta^{\prime}\in\mathbb{R}$,
[…]
Proof.
Let $X\in W^{j,q}_{\delta}$. Lemma A.7 implies
$\|X\|_{W^{j,q}_{\delta}}\lesssim\|PX\|_{W^{j-2,q}_{\delta-2}}+\|X\|_{W^{j-2,q}_{\delta}}$ (A.12)
and we wish to improve the final term from $\|X\|_{W^{j-2,q}_{\delta}}$ to $\|X\|_{W^{j-2,q}_{\delta^{\prime}}}$.
Let $\chi(x)$ be a decreasing smooth function vanishing for $|x|>2$ and satisfying $\chi(x)=1$ for $|x|<1$, and set $\chi_{R}(x):=\chi(x/R)$ with $R>0$ sufficiently large chosen below. Decompose $X=X_{0}+X_{\infty}$, with $X_{0}:=\chi_{R}X$ and $X_{\infty}:=(1-\chi_{R})X$. Since $X_{0}$ is supported in a ball
$\|X\|_{W^{j-2,q}_{\delta}}$ $\leq\|X_{0}\|_{W^{j-2,q}_{\delta}}+\|X_{\infty}\|_{W^{j-2,q}_{\delta}}$
$\lesssim\|X\|_{W^{j-2,q}_{\delta^{\prime}}}+\|X_{\infty}\|_{W^{j,q}_{\delta}}$ (A.13)
where we have also used a trivial estimate for $X_{\infty}$. Inequality (A.9) of Proposition A.8 implies
$\|X_{\infty}\|_{W^{j,q,}_{\delta}}$ $\lesssim\|\overline{P}X_{\infty}\|_{W^{j-2,q}_{\delta-2}}+\|X_{\infty}\|_{W^{j-2,q}_{\delta^{\prime}}}$
$\leq\|PX_{\infty}\|_{W^{j-2,q}_{\delta-2}}+\|(P-\overline{P})X_{\infty}\|_{W^{j-2,q}_{\delta-2}}+\|X_{\infty}\|_{W^{j-2,q}_{\delta^{\prime}}}.$ (A.14)
Observe $(P-\overline{P})X_{\infty}=\chi_{R/2}(P-\overline{P})X_{\infty}$. We can write
$(P-\overline{P})=\sum_{|\alpha|\leq 2}A^{\alpha}\partial_{\alpha}$
with matrix coefficients $A^{\alpha}\in W^{k-2+|\alpha|,p}_{\tau-2+|\alpha|}$ and Lemma A.12 proved below implies $\chi_{R/2}A^{\alpha}\to 0$ in $W^{k-2+|\alpha|,p}_{\tau-2+|\alpha|}$ as $R\to\infty$. Lemma A.3 then implies $\chi_{R/2}(P-\overline{P})$ converges to zero in operator norm on $W^{j,q}_{\delta}$. Hence we can take $R$ sufficiently large to absorb the term $\|(P-\overline{P})X_{\infty}\|_{W^{j-2,q}_{\delta-2}}$ into the left-hand inequality (A.14) to find
$\|X_{\infty}\|_{W^{j,q,}_{\delta}}\lesssim\|PX_{\infty}\|_{W^{j-2,q}_{\delta-2}}+\|X\|_{W^{j-2,q}_{\delta^{\prime}}}.$
Now decompose $PX_{\infty}=\chi_{R}PX+[P,\chi_{R}]X$ and note that the coefficients of $P,\chi_{R}]$ are supported on $B_{R}$. A computation using the bounded domain analog of Lemma A.3 (see holst_scaling_2023 Theorem 2.5) shows $P,\chi_{R}]$ is continuous as a map $W^{j,q}(B_{R})\to W^{j-1,q}(B_{R})$. This fact and the estimate $\|\chi_{R}PX\|_{W^{j-2,q}_{\delta-2}}\lesssim\|PX\|_{W^{j-2,q}_{\delta-2}}$ imply
$\|X_{\infty}\|_{W^{j,q,}_{\delta}}\lesssim\|PX\|_{W^{j-2,q}_{\delta-2}}+\|X\|_{W^{j-1,q}(B_{R})}+\|X\|_{W^{j-2,q}_{\delta^{\prime}}}.$ (A.15)
For each $\epsilon>0$, Sobolev interpolation implies that there is a $C_{\epsilon}>0$ such that
$\|X\|_{W^{j-1,q}(B_{R})}\leq\epsilon\|X\|_{W^{j,q}(B_{R})}+C_{\epsilon}\|X\|_{W^{j-2,q}(B_{R})}\lesssim\epsilon\|X\|_{W^{j,q}_{\delta}}+C_{\epsilon}\|X\|_{W^{j-2,q}_{\delta^{\prime}}}$ (A.16)
where the implicit constant is independent of $\epsilon$. Combining inequalities (A.12)-(A.16) and choosing $\epsilon$ sufficiently small so as to absorb the term $\epsilon\|X\|_{W^{j,q}_{\delta}}$ into the left-hand side we obtain the desired inequality:
$\|X\|_{W^{j,q}_{\delta}}\lesssim\|PX\|_{W^{j-2,q}_{\delta-2}}+\|X\|_{W^{j-2,q}_{\delta^{\prime}}}.$
It remains to prove the following technical lemma used in the proof of Proposition A.11.
Lemma A.12.
Let $\chi$ be an increasing smooth function that equals zero on $B_{1}$ and equals $1$ on $E_{2}$, and define $\chi_{R}(x)=\chi(x/R)$. Given $u\in W^{j,q}_{\delta}$,
[…]
Proof.
Consider a scaling operator $S_{r}$. Then $S_{r}^{*}(\chi_{R}u)$ vanishes on $A_{1}$ unless $r\geq R$. Moreover, it equals $S_{r}^{*}u$ on $A_{1}$ unless $1\leq r/R\leq 2$, in which case the values and derivatives of $S_{r}^{*}(\chi_{R})$ are uniformly bounded independent of $r$ and $R$. Hence for $R\geq 1$,
[…]
Taking the limit in $R$ proves the result.
Combining estimate (A.11) from Proposition A.11 with the compact embedding $W^{j,q}_{\delta}\hookrightarrow W^{j-2,q}_{\delta^{\prime}}$ for $\delta^{\prime}>\delta$, a standard argument (e.g. Theorem 1.10 of bartnik1 ) implies that $P$ is semi-Fredholm:
Corollary A.13.
Suppose that $g_{ab}$ and $N$ satisfy Assumption A.1 with parameters $k$, $p$, and $\tau$. Suppose $(j,q)\in\mathcal{S}^{k,p}$, and that $\delta$ is non-exceptional. Then
$P:W^{j,q}_{\delta}\rightarrow W^{j-2,q}_{\delta-2}$
is semi-Fredholm (that is, it has a finite dimensional kernel and a closed range).
As a final step before establishing the full Fredholm theory, we now show that the kernel of $P$ depends on $\delta$ but not on the Sobolev parameters. This fact is not immediate from Lemma (A.7), which allows for an improvement of the number of derivatives, but only at a fixed integrability parameter $q$.
Lemma A.14.
[…]
Proof.
Suppose $K\in W^{j,q}_{\delta}$ lies in the kernel of $P$. We first show that $K\in W^{j,q}_{\delta^{\prime}}$ for some $\delta^{\prime}<\delta$.
Indeed,
$\overline{P}K=(\overline{P}-P)K+PK=(\overline{P}-P)K.$
The decay of the coefficients of $(\overline{P}-P)$ then imply $\overline{P}K\in W^{j-2,q}_{-2+\delta+\tau}$. Now pick $\delta^{\prime}>\lfloor\delta\rfloor$ with $\delta+\tau<\delta^{\prime}<\delta$. Because the interval $\delta^{\prime},\delta]$ contains no integers, the obstruction conditions of Proposition A.8 part (b) are the same for $\delta$ and $\delta^{\prime}$ and hence there is a $Y\in W^{j,q}_{\delta^{\prime}}$ with $\overline{P}K=\overline{P}Y$. We conclude $Y$ and $K$ differ at most by a polynomial of degree less than $\delta^{\prime}$ and consequently $K\in W^{j,q}_{\delta^{\prime}}$ as well.
Now suppose $j<k$. If $(j+1,q)\in\mathcal{S}^{k,p}$ we can apply Lemma A.7 to conclude $K\in W^{j+1,q}_{\delta}$. Otherwise inequality (A.1) shows that we can lower $q$ to some $Q$ so that $(j+1,Q)\in\mathcal{S}^{k,p}$. Although $W^{k,q}_{\delta}$ does not embed in $W^{k,Q}_{\delta}$ directly, we can use the faster decay rate and the embedding $W^{j,q}_{\delta^{\prime}}\hookrightarrow W^{j,Q}_{\delta}$ along with Lemma A.7 to conclude $K\in W^{j+1,Q}_{\delta}$. After finitely many iterations of this process we conclude $K\in W^{k,Q}_{\delta}$ for some $Q$ with $(k,Q)\in\mathcal{S}^{k,p}$. We now wish to improve $Q$ to $p$.
In fact, from our arguments above we know $K\in W^{k,Q}_{\delta^{\prime}}$ for some $\delta^{\prime}<\delta$. But $W^{k,Q}_{\delta^{\prime}}\hookrightarrow W^{k-1,Q^{\prime}}_{\delta}$ for all $Q^{\prime}<\infty$ with $1/Q^{\prime}\geq 1/Q-1/n$. If 1/Q1/n>1/p1/Q-1/n>1/p then we can take $1/Q^{\prime}=1/Q-1/n$ and apply Lemma A.7 to conclude $K\in W^{k,Q^{\prime}}_{\delta}$. After finitely many iterations we find $K\in W^{k-1,Q^{\prime\prime}}_{\delta^{\prime}}$ with $Q^{\prime\prime}\geq p$ and $\delta^{\prime}<\delta$. But then $K\in W^{k-1,p}_{\delta}$ as well and a last application of Lemma A.7 shows $K\in W^{k,p}_{\delta}$.
Finally, repeating the decay lowering argument from the start of the proof finitely many times depending on the size of $\tau$ we conclude $K\in W^{k,p}_{\delta^{\prime}}$ for all $\lfloor\delta\rfloor<\delta^{\prime}\leq\delta$.
We now obtain the primary Fredholm result.
Proposition A.15.
Suppose that $g_{ab}$ and $N$ satisfy Assumption A.1 with parameters $k$, $p$, and $\tau$. Suppose $(j,q)\in\mathcal{S}^{k,p}$ and $\delta$ is non-exceptional. Then
$P:W^{j,q}_{\delta}\rightarrow W^{j-2,q}_{\delta-2}$
is Fredholm, and its Fredholm index satisfies $\iota(P^{j,q}_{\delta})=\iota(\overline{P}^{j,q}_{\delta})$ and is independent of $j$ and $q$.
Let $P^{*}$ denote the action of $P$ on $W^{2-j,q^{\prime}}_{2-n-\delta}$. Then the kernel of $P^{*}$ is independent of $j$ and $q$ and depends only on $\delta$. Given $V\in W^{j-2,q}_{\delta-2}$, the equation
$PX=V$
is solvable for $X\in W^{j,q}_{\delta}$ if and only if $\int_{\mathbb{R}^{n}}\left<V,K\right>_{g}dV_{g}=0$ for all vector fields $K\in\ker(P^{*})$.
Proof.
Corollary A.13 implies that $P$ is semi-Fredholm. Lemma A.6 shows that the adjoint of $P$ is $P:W^{2-j,q^{\prime}}_{2-n-\delta}\to W^{-j,q^{\prime}}_{-n-\delta}$, which we denote by $P^{*}$. Since $2-n-\delta$ is nonexceptional when $\delta$ is, Corollary A.13 implies that $P^{*}$ is also semi-Fredholm. We claim the following:
$\mathrm{im}(P)=(\ker(P^{*}))^{\perp}.$ (A.17)
Together with Lemma A.5, this establishes the claimed solvability criterion. Since the kernel of $P^{*}$ is finite dimensional, this also establishes that $P$ is Fredholm. Lemma A.14 shows that the kernel of $P^{*}$ is independent of $j$ and $q$.
Lemma A.6 implies
$\int_{\mathbb{R}^{n}}\left<PX,Y\right>_{g}dV_{g}=\int_{\mathbb{R}^{n}}\left<X,P^{*}Y\right>_{g}dV_{g}$
for all $X\in W^{j,q}_{\delta}$ and $Y\in W^{2-j,q^{\prime}}_{2-n-\delta}$, and the inclusion $\mathrm{im}\,P\subset(\ker P^{*})^{\perp}$ is immediate. Conversely, suppose that $Z\in W^{j-2,q}_{\delta-2}$ is not in the image of $P$. Since $P$ is semi-Fredholm its image is closed and the Hahn-Banach Theorem together with Lemma A.5 imply that there exists a $Y\in W^{2-j,q^{\prime}}_{2-n-\delta}$ satisfying both
$\int_{\mathbb{R}^{n}}\left<PX,Y\right>_{g}\;dV_{g}=0$ (A.18)
for all $X\in W^{j,q}_{\delta}$, and
$\int_{\mathbb{R}^{n}}\left<Z,Y\right>_{g}\;dV_{g}\neq 0.$ (A.19)
Equation (A.18) implies that
$\int_{\mathbb{R}^{n}}\left<X,P^{*}Y\right>_{g}\;dV_{g}=0$
for all $X\in W^{j,q}_{\delta}$, so Lemma A.5 further yields that $Y\in\ker P^{*}$. That is: If $Z$ is not in the image of $P$, then there exists $Y\in\ker(P^{*})$ satisfying equation (A.19), and hence $Z\not\in(\ker(P^{*}))^{\perp}$. This confirms equation (A.17).
Regarding the index, the paths of metrics $g_{t}=(1-t)\overline{g}+tg$ and lapses $N_{t}=(1-t)+tN$ yield a continuous path of operators $P_{g_{t},N_{t}}$ from $\overline{P}$ to $P_{g,N}$, and local constancy of the index shows that $\iota(P^{j,q}_{\delta})=\iota(\overline{P}^{j,q}_{\delta})$. The fact that the index is independent of $(j,q)$ follows from the corresponding fact for $\overline{P}$ from Proposition A.8.